A Complete Bibliography of Philosophy of Science: 1970–1979

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

04 March 2019 Version 1.03

Title word cross-reference

F [460].

1950 [358]. 19th [654]. 19th-Century [654].

A. [62]. Aage [14]. Aaron [713]. Absolute [406, 533]. Abstracts [55, 37, 66, 127, 144, 84, 194, 171, 264, 380, 399, 409, 488]. Absurdity [680, 681]. Accelerated [295]. Acceptance [102, 101]. Account [582, 376]. Achenstein [167]. Achinstein [697, 424, 77, 124]. Ackermann [630]. Ackoff [486]. Action [322]. Acts [665]. Adams [588]. Adaptation [90, 227, 226, 63]. Add [329]. Additive [22]. Adequacy [217, 115]. Adolf [230, 425]. Advice [626]. Against [121, 79, 217]. Ager [487]. Aim [357]. Alan [162, 211]. Alasdair [395]. Alastair [338]. Alexander [569]. Alfred [287]. Algebraic [584]. Alleged [642]. Alternative [582]. Alternatives [132]. Alwyn [287]. American [664]. among [322]. Analogies [181, 136, 77, 182]. Analyses [440]. Analysis [96, 330, 516, 78, 150, 655, 394, 459, 275, 549, 222, 569, 53]. Analytic [493]. Analytische [82, 208]. Angel [103]. Angelelli [106]. Announcement [471, 557]. Anomalies [141]. Answers [629]. Application [376]. Applied [378]. Appraisal [568, 604]. Appraisals [626]. Approach [54, 465, 474, 584, 176, 236]. Approximations [640]. Archaeological [701]. Archeology [236]. Archimedean [367]. Argument [79, 680, 681, 384]. Arguments [217, 646, 116]. Aristotelean [149]. Aristotelian [519]. Arnheim [262]. Arnold [64]. Arrow [657]. Art [54, 94, 280, 467]. Articles [305]. Artificial [647]. Arturo [238]. Aspects [639, 99, 280, 405, 32]. Assessment [563]. Assimilation [660]. Assumptions [574, 6, 23]. Asymmetries [319]. At-At [529]. Atomism [654, 549]. Attempt [329, 260]. Attention [260]. Attributions [690]. Audi [484]. Austin [12]. Authors [260]. Autonomy [561]. Axiomatic [122]. Axiomatization [2, 675]. Axioms [403, 367].

B [585, 629, 663, 394]. Back [18, 38, 56, 68, 85, 110, 128, 146, 172, 195, 213, 242, 265, 290, 315, 326, 340, 353, 361, 382, 401, 410, 432, 444, 455, 472, 489, 508, 523, 544, 558, 571, 592, 610, 632, 652, 667, 688, 704, 719, 421]. Backward [522, 457]. Band [82]. Banff [309]. Barker [167]. Barrett [567]. Barry [703]. Bas [335]. Basic [250]. Basis [469]. Basri [139]. Bastin [233]. Be [221]. Beauchamp [278]. Becoming [319]. Been [515]. Begrundung [82]. Behavior [532, 459, 365, 396]. Behavioral [669]. Behaviorism [459]. Being [421, 244]. Beliefs [635]. Belnap [629]. Berg [80]. Bergmann [505]. Bernard [260]. Bertalanffy [52]. Bertrand [427]. Beth [40]. Better [670, 386, 501]. between [221, 208, 253, 161, 60, 279, 369]. Beyond [94, 233]. Bias [414]. Bifurcation [437]. Biological [206, 198, 710, 9, 90, 227, 614, 226, 205]. Biology [1, 440, 219, 693, 246, 438, 76, 566, 35, 405, 372, 336, 168]. Blackburn [339]. Blalock [240]. Blokhintsev [13]. Blume [684]. Boden [396, 647]. Bogdan [521]. Boltzmann [585]. Bolzano [260]. Bonelli [607]. Book [230, 16, 521, 187, 628, 665, 34, 283, 337, 351, 287, 231, 54, 664, 372, 310, 312, 210, 374, 262, 189, 313, 426, 307, 314, 139, 167, 13, 14, 379, 188, 485, 606, 470, 588, 233, 352, 163, 166, 15, 158, 164, 169, 258, 284, 305, 468, 430, 585, 687, 686, 647, 609, 424, 630, 81, 506, 591, 648, 106, 107, 358, 700, 586, 568, 304, 282, 683, 649, 257, 398, 425, 699, 397, 567, 336, 629, 308, 393, 309, 82, 260, 466, 375, 229, 303, 663, 604, 162, 192, 587, 714, 605, 650, 542, 484, 288]. Book [569, 394, 685, 108, 289, 467, 239, 235, 335, 703, 584, 212, 259, 607, 711, 311, 52, 140, 157, 159, 191, 234, 237, 261, 360, 684, 306, 286, 338, 339, 376, 717, 715, 373, 702, 238, 666, 427, 190, 209, 285, 395, 211, 240, 378, 543, 141, 161, 168, 428, 469, 662, 701, 125, 590, 713, 165, 334, 608, 396, 486, 64, 541, 716, 143, 631, 371, 236, 377, 32, 33, 35, 487, 520, 53, 232, 429, 63, 105, 589, 160]. Books [17, 36, 65, 83, 109, 126, 142, 170, 193, 263, 324, 359, 400, 431]. Boston [125]. Bradie [453]. Bradley [188]. Brain [328, 238]. Brains [417].

Bridge [597]. British [127]. Brittan [235, 715]. Brody [296]. Brownian [530]. Bruce [701]. Bruno [303]. Bryan [375]. Bub [542]. Buchdahl [189]. Building [137]. Bunge [393, 311, 155, 377]. Burn [643]. Butts [702, 33, 254]. Bynum [305]. Byrne [143].

C [379, 591, 229, 335, 584, 62, 702, 600, 141, 161, 168, 469, 63, 536]. C-Fiber [536]. Cahn [81]. Cancellation [497]. Carl [32]. Carnap [229, 493, 512]. Case [522, 457, 132, 636]. Cassirer [310]. Causal [277, 152, 385, 624, 461, 529, 528, 645, 644, 642, 176]. Causal-Explanatory [461]. Causality [278, 465, 641, 234, 285]. Causation [522, 498, 207, 457, 461, 517, 565, 426, 157]. Cause [221]. Causes [255, 656, 254]. Caws [107]. Cement [426]. Century [351, 687, 334, 654]. Challenge [352]. Chance [261]. Change [694, 16, 469]. Charles [567, 236, 429]. Chihara [429]. Chissick [686]. Choosing [60]. Chronometry [230]. Circle [429, 662]. Clarke [681]. Classes [495]. Classical [522, 457, 660, 699, 659, 285]. Claudio [487]. Clendinnen [698]. Clock [138, 537]. Cognitive [94, 280, 150]. Cohen [683, 125, 237]. Coherence [360, 294]. Colin [604]. Collected [567, 160]. Colodny [352]. Combined [138]. Comment [556, 677, 681, 222]. Comments [440, 554, 228]. Common [590]. Community [428]. Comparability [130]. Components [477]. Computation [344]. Computer [713]. Computers [322]. Comte [72]. Conceivability [75, 682]. Concept [310, 562, 137, 288, 202, 439, 376, 116]. Conception [301, 525]. Concepts [185, 714, 614, 124, 485]. Conceptual [379, 556, 555, 305]. concerning [496]. Conclusions [638]. Conditional [350, 464]. Conditionalization [479, 613]. Conditionals [494, 5, 588]. Conditions [115, 246]. Conduct [397]. Conference [309]. Confirmation [616, 123, 514, 42, 115, 4, 658, 600, 60, 441]. Confirmed [513]. Conjoint [22]. Connections [44]. Consciousness [322, 505]. Conservation [514]. Considerations [118]. Consistency [97, 98]. Constance [169]. Constancy [535]. Constant [260]. Construction [240]. Contemporary [379, 185]. Contexts [461, 642]. Contextual [218]. Continuous [478, 552]. Contraction [138]. Contraction-Clock [138]. Contradictories [497]. Contraries [497]. Contributions [64]. Control [396]. Controversy [639]. Convenience [510]. Convention [201]. Conventionalism [24]. Conventionality [491, 368]. Copernican [391]. Copi [284, 464]. Cornman [590]. Corrections [522]. Correspond [573]. Correspondence [88, 617, 406]. Cosmolog [363]. Cosmology [97, 98]. Counter [463, 225]. Counter-Intuitive [463, 225]. Counterfactuals [430]. Countering [225]. Counting [153, 45]. Covering [135, 298]. Cracking [533]. Creation [664]. Crisis [549]. Criteria [435]. Criterion [343, 600, 676]. Critical [516, 58, 102, 390, 536, 543, 716]. Criticism [116, 162]. Critics [61, 428]. Critique [374, 282, 699, 174, 63]. Cross [12]. Crossley [468]. Crossroads [283]. Cultural [662]. Current [549, 63]. Curvature [696]. Cybernetics [650].

D [13, 700, 257, 629, 261, 496, 46, 346]. D-N [496, 46, 346]. Darwin [567, 428, 662, 132, 428]. Darwinian [312]. Data [516, 415]. David [430, 663, 428, 255, 98]. Debates [706]. Decision [700]. Decline [92]. Deductive [139, 480, 47, 596]. Deductive-Nomological [596]. Defeasibility [599]. Defence [338]. Defended [364]. Defense [641, 296]. Definition [493, 617, 202, 407]. Degeneration [390]. Degrees [200]. del [608]. Delaware [187]. Demarcation [343]. Demon [344]. Demonstrative [293]. Dependence [88]. Depth [691]. Derek [487]. Describing [177]. Description [299]. Description-Relativity [299]. Desmond [681]. Detailed [260]. Determination [656]. Determinism [215, 583, 504, 178]. Development [89, 550, 427]. Deviant [586]. Dewey [543]. Dichotomous [516]. Diesing [259]. Difference [221, 279]. Dimension [692]. Dimensionally [87, 573]. Dingle [283]. Dirac [522, 457]. Direct [207]. Direction [329]. Disconfirmation [224]. Discovery [180, 151, 579, 342, 357, 259, 317]. Discussion [58, 102, 208]. Disease [565]. Dispositions [248]. Disprove [623]. Dissonance [150]. Distant [676]. Distinction [119, 253, 369]. Divine [35]. Do [392]. Doctrines [374]. Does [623, 281, 317]. Dogma [581]. Dogmas [503]. Domain [231, 352]. Dominance [206]. Dominant [116]. Don [281]. Doubts [646, 415]. Drawing [119]. Dretske [277, 301, 618]. Dualism [92]. Duhem [419, 621]. Duhemian [388].

Earlier [260]. Early [714, 285]. Earth [487]. Eberle [232]. Econometric [25]. Economic [24, 78, 210, 374, 699]. Economics [374, 568, 699]. Economy [449]. Editorial [560, 634]. Edmond [628]. Edmund [143]. Edward [470, 288, 662]. Edwards [140, 376]. Effect [221]. Einstein [371, 156, 151, 93]. Electrodynamics [522, 574, 457, 660, 659, 522]. Eliminating [709]. Elimination [197]. Eliminative [99, 709]. Elkana [683]. Elucidation [645]. Emendation [621]. Emergence [185, 314]. Emery [486]. Empirical [697, 217, 583, 209]. Empiricist [420]. Encountering [463]. Encyclopedia [140]. Energy [70, 549]. England [351, 334]. Englehardt [685]. Entails [120]. Entities [499, 272]. Entity [565]. Entrenchment [625]. Entropy [210]. Epistemic [547, 527]. Epistemics [374]. Epistemological [268, 683]. Epistemologizing [636]. Epistemology [450]. Equivalence [253, 370]. Erfahrung [208, 466]. Ergodic [454, 575, 269]. Erhard [394]. Erklarung [82]. Ernest [16, 588, 228]. Ernst [64, 310, 125]. Erwin [165, 288]. Essay [424, 117]. Essays [287, 467, 239, 701, 32]. Essential [648]. Essentialism [296]. Ethics [685]. Ethnosemantics [477]. Euclidean [180]. Evaluating [177, 482]. Evaluation [28, 547]. Event [331]. Events [112, 665]. Evidence [26, 655, 27]. Evidential [414, 655]. ’Evolution [578, 556, 555, 428, 584]. Evolutionary [1, 89, 469, 116, 320, 63]. Examination [516, 11]. Exercise [78]. Expectancy [692]. Expected [479]. Experience [208]. Experiment [249, 607]. Experimental [313, 459]. Experiments [547]. Explanation [71, 149, 532, 548, 712, 582, 424, 435, 496, 416, 220, 714, 519, 184, 615, 47, 154, 252, 135, 299, 46, 575, 596, 285, 459, 525, 517, 396, 269, 346, 205, 32, 674, 179, 163, 609, 192, 236]. Explanations [1, 582, 690, 480, 452]. Explanatory [637, 461, 691, 575, 642, 714]. Explication [434]. Explicitly [236]. Exploration [416]. Exposition [58, 260]. Extension [40]. Extensionality [461, 642]. Extensive [601, 366]. Ey [165].

F [34, 606, 683, 685, 376, 161, 143]. Fact [239]. Facts [71, 591, 530]. Fall [92]. Fallibilism [390]. False [492]. Falsifiability [58]. Falsifiable [320]. Falsification [218, 274]. Fate [81]. Feature [131]. Feibleman [313]. Festinger [150]. Feyerabend [530, 390]. Fiber [536]. Fictionalism [197]. Fiery [74]. Fifty [686]. Finetti [303]. Finite [217, 675]. Finocchiaro [609]. Finzi [487]. First [309, 318]. Fisico [608]. Fit [675]. Fitness [673]. Fitzgerald [138]. Five [388]. Flagpoles [580]. Flow [475]. Fly [281]. Fodor [163, 253, 369]. Force [318]. Forces [292]. Form [661]. Formal [77, 660]. Formation [137, 202]. Formulations [240]. Foulkes [585]. Foundations [187, 268, 379, 245, 686, 198, 308, 311, 158]. Four [20, 543]. Fraassen [348, 335]. France [43]. Francia [608]. Fraser [605]. Fred [486]. Frederick [687, 398]. Frege [305, 106, 190]. Friedlander [397]. Friedrich [190]. Friedrichs [212]. Front [19, 39, 57, 69, 86, 111, 129, 147, 173, 196, 214, 243, 266, 291, 316, 327, 341, 354, 362, 383, 402, 411, 433, 445, 456, 473, 490, 509, 524, 545, 559, 572, 593, 611, 633, 653, 668, 689, 705, 721]. Function [546, 690, 222]. Functional [440, 690, 95, 370, 76, 275]. Functionalism [157, 707]. Functions [575]. Fundamental [318]. Furnace [74].

G [374, 352, 282, 235, 717, 715, 32]. Galileo [702]. Ganz [307]. Gaukroger [714]. Geach [373]. Gellner [16]. General [304, 682]. Generality [661]. Generalization [389]. Generis [491]. Genetic [469]. Genetics [577, 219, 679, 693]. Genidentity [385]. Gentzen [160]. Geology [35, 663]. Geometry [180, 695, 230]. George [666, 63, 260, 535]. Georgescu [210]. Georgescu-Roegen [210]. Gerald [371]. Gerd [189]. Gerhard [160]. Germany [687]. Ghiselin [312]. Giannoni [678]. Gibson [636]. Girill [181]. Giuliano [608]. Given [131]. Global [404]. Gödel [59]. Goodman [467, 54, 616]. Gordon [235]. Gottlob [305, 106, 190]. Gould [649]. Grades [414]. Grain [708]. Grammar [217]. Graves [379]. Gravitation [446]. Gravitational [292]. Great [536]. Gregory [687]. Grene [680]. Group [639, 116]. Growth [598, 162]. Grue [476]. Gruesome [481]. Grunbaum [659, 425, 230]. Gustav [505].

H [567, 685, 261, 168, 64]. Haack [586]. Halbband [466]. Hall [166]. Hand [15]. Hannay [338]. Hans [190]. Harre [498, 391, 108]. Hasty [636]. Hatcher [158]. Hawkins [98]. Health [562, 202]. Helmholtz [683].

Hempel [61, 582, 534, 525, 32]. Henri [165]. Herbert [283]. Hermann [683]. Hermes [190]. Hidden [552]. Hiddenness [530]. Hide [337]. High [549]. Higher [294]. Hilary [258, 587]. Hilbert [450, 169]. Hilpinen [101]. Hindess [703]. Hintikka [191]. Historical [423, 239, 584, 12, 541]. History [510, 293, 358, 487, 609, 378]. Holder [403]. Hollis [699]. Holton [371]. Homogeneity [691]. Honor [287, 467]. Hooker [584]. Hooykaas [35]. Howson [604]. Hubert [240]. Hull [428]. Human [304]. Humanistic [118]. Hume [74]. Humphreys [141]. Hylas [15]. Hypotheses [28, 321]. Hypothesis [664, 198, 655, 535, 138]. Hypothetico [313]. Hypothetico-Experimental [313].

I-Counting [153]. Iculties [223, 3]. Ideal [598, 249]. Idealism [273]. Idealizations [96, 330, 640]. Ideas [166]. Identity [623, 492, 198, 331]. If [223, 3]. If-Iculties [223, 3]. Ignacio [106]. Ignorance [364]. II [605, 357, 23]. III [365]. Illness [202]. Images [338]. Immodest [73, 204]. Immodesty [332, 347]. Implication [442]. Implications [237]. Impossibility [267]. Imre [162, 159]. Incommensurable [672]. Incommensurables [132]. Independence [540]. Independent [602, 403]. Indeterminacy [437]. Indeterminism [527]. Indices [102]. Indiscernibles [623, 492]. Individualism [615]. Individualistic [495]. Individuality [612]. Individuation [554, 553]. Induction [293, 48, 698, 303, 540, 716, 521, 237]. Inductive [102, 229, 73, 332, 101, 159, 347, 525, 204, 176]. Inductive-Statistical [525]. Inertia [292, 446]. Inextricability [510]. Inference [413, 102, 525, 191, 376, 700]. Infinite [256, 387, 500]. Influence [529]. Information [67, 145, 241, 325, 381, 443, 507, 570, 651, 720, 28, 637, 191]. Inquiry [55, 144, 194, 264, 380, 399, 409, 488, 274, 231]. Instances [224, 600]. Instrumentalism [654]. Integrative [185]. Intelligence [647]. Intensional [620]. Intensionality [298]. Intentionality [21]. Intentions [134]. Interested [255]. Interface [187]. Interfield [511]. Interpretation [136, 423, 364, 542, 484, 673, 200, 701]. Interpretations [216, 541]. Intrinsic [696, 478]. Introduction [560, 304, 235, 335, 543]. Introspections [412]. Intuitive [463, 225]. Invariance [323]. Invariant [87, 573]. Involving [136]. Irregularity [277, 152, 624]. Irving [284]. Isajiw [157]. Ishiguro [337]. Israel [467, 543]. Issues [560, 563, 336, 586]. Itzkoff [310].

J [521, 426, 379, 188, 605, 394, 191, 286, 702, 125, 53]. Jackson [49, 698, 676]. Jager [427]. James [313, 543, 590, 114]. Jammer [541]. Jan [80]. Janet [206]. Jarvis [665]. Jay [606, 649]. Jeffrey [542, 613, 229]. Jerome [306]. Jerry [163]. Jo [236]. Joan [307]. Jochen [64]. John [630]. Joseph [34, 257, 309]. Journal [127]. Juan [281]. Judging [435]. Judith [665]. Justice [564]. Justification [419, 579].

Kambartel [190]. Kant [333, 715]. Karel [235]. Karl [630, 566, 375].

Kaulbach [190]. Kelvin [132]. Kenneth [650]. Kepler [371]. Kerr [502]. Kim [47]. Kinds [458, 710]. King [43]. Kitts [663]. Knowing [286]. Knowledge [310, 301, 91, 598, 306, 162]. Kordig [300]. Krips [121]. Kuhn [648, 345, 203, 483]. Kuratowski [105]. Kurt [64].

L [664, 374, 426, 206, 607, 486, 236]. L.J [237]. Laboratory [313]. Ladenness [300]. Lakatos [162, 159]. Lambert [235]. Landé [287]. Language [587, 337]. Languages [113, 54]. Laplace [664]. Latsis [568]. Laudan [711]. Lavoisier [643]. Law [664, 210, 292, 424, 135, 298, 35, 424]. Lawlikeness [102, 347]. Lawrence [520, 605]. Laws [215, 582, 531, 594, 597, 318, 11, 87, 573, 569, 618, 228, 12]. Learning [412]. Leblanc [236]. Legacy [167]. Lehman [9, 297]. Leibniz [492, 337]. Lending [15]. Lewis [430, 332, 204]. Lewontin [469]. Life [166]. Likelihood [376, 376]. Limitations [246]. Limits [231]. L’Indagine [608]. Linguistic [328, 606]. Little [329]. Local [521]. Logic [628, 337, 451, 588, 458, 468, 460, 81, 133, 586, 629, 260, 229, 101, 199, 209, 317, 355, 581, 176, 164, 258, 159, 395, 64, 467, 373]. Logical [671, 167, 284, 257, 394, 384]. Logico [328, 584]. Logico-Algebraic [584]. Logico-Linguistic [328]. Lorentz [138]. Lowe [683]. Ludwig [52, 585].

M [284, 609, 81, 681, 239, 607, 240, 160]. M. [680]. Mach [125, 188]. Machamer [717]. Machan [422]. Machine [646, 253, 370]. Mackie [426]. MacKinnon [518]. Magee [375]. Magic [271]. Main [260]. Making [246]. Malcolm [683]. Man [310, 647, 699, 336]. Manier [662]. Manifolds [478]. Many [501]. Margaret [647, 396]. Margolis [418]. Marian [209]. Mario [393, 311, 377]. ’Marketing [408]. Markus [289]. Marriage [510]. Marshall [308]. Masking [645]. Materialism [687, 99, 709]. Materialist [268]. Mathematical [468, 257, 240, 64]. Mathematics [158, 587]. Matter [18, 38, 56, 68, 19, 39, 57, 69, 85, 110, 128, 146, 86, 111, 129, 147, 172, 195, 213, 242, 173, 196, 214, 243, 265, 290, 315, 326, 266, 291, 316, 327, 340, 353, 361, 382, 341, 354, 362, 383, 401, 410, 432, 444, 402, 411, 433, 445, 455, 472, 489, 508, 456, 473, 490, 509, 523, 544, 558, 571, 524, 545, 559, 572, 592, 610, 632, 652, 593, 611, 633, 653, 667, 688, 704, 719, 668, 689, 705, 721, 697, 612, 587, 261, 603, 166, 377]. Matters [373]. Maurice [165]. Max [282, 541]. Maximal [29]. Maxwell [344]. McGuinness [585]. McShane [314]. Mead [543]. Meaning [408, 477, 292, 694]. Meaning-Change [694]. Meaningful [573]. Meaninglessness [288]. Meanings [274]. Measure [637]. Measurement [245, 601, 366, 318, 22, 367, 228]. Mechanical [447, 178, 552]. Mechanics [623, 348, 686, 574, 216, 660, 121, 584, 442, 541, 13, 308, 542, 484, 394]. Mechanism [680, 681, 75, 205]. Mechanisms [365]. Medical [685]. Medicine [560, 565, 628]. Medieval [285, 143]. Meiland [483]. Mellor [261]. Memorium [718]. Memory [384]. Men [52]. Meno [349, 453]. Mental [202, 338]. Merton [334]. Merwe [287]. Metamathematics [104]. Metamethodological [692]. Metaphysical [53]. Metaphysics [446, 189].

Metatheoretical [512]. Method [313, 122, 700, 587, 174, 312, 239, 666, 716, 33, 568, 604, 377]. ’Méthode [72]. Methodological [118, 615, 279, 308, 631]. Methodology [78, 503, 218, 302, 703]. Methods [73]. Metrics [478]. Michael [372, 312, 397, 484]. Michalos [62]. Micro [499, 272, 627]. Micro-Entities [499, 272]. Micro-Particles [627]. Microeconomic [670, 569]. Mind [268, 328, 104, 587, 650, 713, 238]. Mind-Brain [328]. Minded [51]. Minds [417, 12, 52]. Miracle [35]. Mistakes [709]. Modal [348]. Modalities [176]. Model [31, 556, 348, 598, 416, 519, 184, 46, 596, 555, 442, 644, 298, 346, 377, 565]. Models [181, 89, 333, 136, 77, 182, 154, 528, 713, 204, 25]. Modes [8]. Molecular [577, 693, 405]. Mondo [608]. Moor [712]. Moral [561]. Mostowski [105]. Motion [318, 530]. Mr. [518, 9]. Müller [289]. Müller-Markus [289]. Mundle [591]. Munson [336, 10]. Murphy [628]. Musgrave [162, 626]. Must [221]. Mysticism [607]. Myth [522, 457, 536].

N [468, 605, 496, 46, 346]. Nachgelassene [190]. Nagel [228]. Natanson [165]. Natural [664, 556, 458, 647, 710, 555, 45, 389, 279, 35, 63]. Nature [531, 292, 44, 336, 685, 618, 512, 487]. Nebular [664]. Necessary [44]. Negative [448]. Nell [699]. Nelson [54, 467]. Neo [699]. Neo-Classical [699]. Neutral [95]. Newton [292, 318]. Newtonian [276, 660]. Nicholas [210, 358, 192, 360, 395, 716, 631]. Nickles [298]. Niiniluoto [619]. Nineteenth [687]. Nissim [660]. Nissim-Sabat [660]. Nomic [656]. Nominalistic [232]. Nomological [596, 661]. Non [180]. Non-Euclidean [180]. None [501]. Notation [305]. Note [276, 613, 423, 505, 407, 536, 482]. Notes [465]. Notion [384, 389]. Novel [260]. Nuel [629]. Numbers [45, 664]. Numerical [573]. Nut [533].

O [470]. Oaklander [554]. Objection [708]. Objective [475]. Objectivity [48, 302, 345, 483]. Observables [552]. ’Observation [407, 300, 119]. Observational [323]. Observationality [419]. Observing [120]. Omer [184, 252]. One [131, 574, 580, 6, 23, 412]. One-Way [574, 6, 23]. Ontogeny [649]. Ontology [8, 429]. Opacity [175]. Operant [459]. Opinion [143]. Order [366, 294]. Orderings [22]. Organisms [538]. Oriented [357]. Origins [371]. Orman [164]. Orthodox [42]. Other [665, 12, 32].

P [585, 191, 234, 373]. Pair [22]. Panel [58]. Pantin [161]. Papers [567, 160, 587, 587]. Paradigm [549]. Paradigms [352]. Paradox [616, 123, 42, 534, 537, 183, 692, 349, 156, 93, 453, 186]. Paradoxes [352, 616, 121, 600]. Paradoxical [580, 59]. Pargetter [676]. Part [304, 342, 357, 6, 23]. Partial [454]. Particles [522, 457, 627]. Particular [71, 112]. Particulars [554, 250, 553]. Past [576]. Pastore [358]. Patterns [41, 259]. Patty [236]. Paul [567, 259, 140]. Peirce [543, 449, 716]. Perception [301, 358, 7, 114, 636, 396, 717, 591, 590]. Personality [34].

Personnelle [589]. Persons [417]. Perspective [230, 541]. Perspectives [702, 287, 684]. Peter [167, 107, 717]. Petersen [14]. Philip [314]. Philosopher [125]. Philosophical [230, 14, 563, 352, 639, 585, 586, 699, 336, 587, 569, 20, 575, 378, 405, 425, 685]. Philosophie [82, 208]. Philosophy [187, 127, 34, 337, 372, 189, 13, 188, 510, 31, 424, 630, 58, 282, 219, 650, 104, 235, 335, 390, 117, 512, 238, 211, 566, 125, 713, 541, 716, 32, 106, 714, 140, 427, 165, 164, 258, 393, 703]. Phylogeny [649]. Physical [21, 313, 604, 696, 70, 2, 40]. Physicist [125]. Physics [187, 330, 14, 585, 293, 582, 122, 201, 714, 640, 257, 393, 311]. Picturability [272, 627]. Picturable [499]. Pitt [702]. Place [71]. Podolski [156]. Podolsky [93]. Points [25]. Polanyi [349, 453]. Popper [630, 375, 540, 174, 566]. Positional [554, 553]. Positive [26, 224, 403, 72]. Positivism [167]. Possible [452]. Post [143]. Post-medieval [143]. Postan [239]. Postscript [203]. Postulates [550, 88, 62]. Power [637]. Powers [396]. Pragmatic [330]. Pragmatism [631]. Pragmatists [543]. Preacceleration [522, 457]. Predictability [215]. Prediction [435, 246, 658, 339]. Prediction-Making [246]. Predictions [251, 462, 320]. Preference [62]. Preservation [131]. Presuppositions [143]. Price [686]. Primes [186]. Principle [623, 697, 215, 492, 686, 497, 406, 482, 603, 35, 429]. Principles [108]. Prior [321]. Probabilistic [413, 465, 234, 596, 641, 176]. Probabilities [448, 526, 321, 294]. Probability [448, 671, 350, 460, 91, 463, 464, 216, 225, 271, 178, 5, 112, 229, 143, 303, 143]. Problem [328, 350, 329, 464, 342, 159, 552]. Probleme [208, 82]. Problems [585, 425, 7, 202, 20, 711, 306, 311]. Procedure [313]. Process [210]. Professor [554, 418]. Programs [532, 598, 134]. Progress [206, 343, 711]. Progressivism [116]. Projectibility [199]. Projecting [476]. Projection [183]. Proof [267]. Propensity [673]. Proposed [328, 154]. Protophysik [289]. Provincial [438]. Przelecki [209]. Psychiatry [165]. Psychic [70]. Psycholinguists [20]. Psychological [117, 163]. Psychology [309, 113, 286, 441, 309, 396]. Psychoneural [198]. Puccetti [417]. Pure [583]. Purely [413]. Purpose [322, 30]. Purposeful [486]. Purposive [459, 365, 396]. Putnam [587, 100, 258, 587]. Puzzle [496].

Qualitative [573]. Qualities [554, 553]. Quantum [187, 287, 623, 671, 13, 352, 348, 363, 686, 133, 122, 216, 121, 92, 542, 484, 394, 584, 178, 541, 552, 14, 233]. Quantum-Mechanical [178]. Quartercentenary [346]. Query [625]. Questions [629]. Quine [164]. Quinn [621].

R [398, 309, 108, 607, 306, 286, 702, 469, 35, 416]. Radical [356]. Radu [521]. Raimo [485]. Ramsey [387, 500]. Ramsey-Reducts [500]. Randomness [314]. Rational [62, 699]. Rationalism [390]. Rationality [561, 342, 357, 27]. Ravetz [306]. Raymond [125]. Re [463]. Re-Encountering [463]. Realism [94, 654, 207, 706, 8, 620]. Realist [245].

Reality [587, 187]. Really [281, 133]. Reason [607, 339]. Reasonable [526, 421]. Reasons [91]. Reassessment [72]. Received [148]. Reception [428]. Reciprocity [247]. Recognition [447]. Record [487]. Redman [236]. Reduction [79, 434, 41, 577, 219, 679, 693, 80]. Reductionism [707]. Reducts [500]. Redundancy [480]. Reexamination [624]. Reference [620, 669]. Referential [175]. Reflexive [251, 462]. Refutation [93]. Refuting [530]. Reichenbach [11]. Reid [169, 180, 535]. Reinterpretation [391]. Rejecting [321]. Rejoinder [104, 622, 539]. Related [305]. Relatedness [645]. Relation [366]. Relations [247, 655, 573, 368, 161, 528, 644]. Relationship [208]. Relative [677, 551, 406, 389]. Relativistic [574, 308]. Relativity [379, 677, 299, 295, 151, 551, 695, 368, 504, 682, 6, 23]. Relativization [527]. Relevance [480, 239, 600, 517]. Religion [351, 666]. Remark [61]. Remarks [137, 186]. Replacement [434]. Replacing [131]. Replies [61]. Reply [181, 277, 255, 502, 278, 619, 206, 58, 660, 518, 102, 48, 50, 678, 10, 227, 627, 535, 418, 275, 297, 80, 98]. Report [412]. Reporting [177]. Representation [518, 606]. Rescher [395, 192, 360, 716, 631]. Research [598, 449]. Resolution [42, 121, 534]. Respect [22]. Response [49, 50, 422]. Resultate [82, 208]. Retardation [138]. Retrocausation [522, 457, 660, 659]. Review [230, 16, 521, 187, 628, 665, 34, 283, 337, 351, 287, 231, 54, 664, 372, 310, 312, 210, 374, 262, 189, 313, 426, 307, 314, 139, 167, 13, 14, 379, 188, 485, 606, 470, 588, 233, 352, 163, 166, 15, 158, 164, 169, 258, 284, 305, 468, 430, 585, 687, 686, 647, 609, 424, 630, 81, 506, 591, 648, 106, 107, 358, 700, 586, 568, 304, 282, 683, 649, 257, 398, 425, 699, 397, 567, 336, 629, 308, 393, 309, 82, 208, 260, 466, 375, 229, 303, 663, 604, 162, 192, 714, 605, 650, 542, 484, 288]. Review [569, 394, 685, 108, 289, 467, 239, 235, 335, 703, 584, 212, 259, 607, 711, 311, 52, 140, 157, 159, 191, 234, 237, 261, 360, 684, 306, 286, 338, 339, 376, 717, 715, 373, 702, 238, 666, 427, 190, 209, 285, 395, 211, 240, 378, 543, 141, 161, 168, 428, 469, 662, 701, 125, 590, 713, 165, 334, 608, 486, 64, 541, 716, 143, 631, 371, 236, 377, 32, 33, 35, 520, 53, 232, 429, 63, 105, 589, 160]. Reviews [587, 396, 487]. Revisited [182, 248]. Revolution [100, 391, 713, 203, 607]. Revolutionary [390]. Richard [351, 231, 229, 467, 718]. Riddle [356]. Righini [607]. Rival [672, 482]. Robert [352, 630, 683, 212, 717, 125, 334, 33]. Robots [52]. Rod [138]. Roegen [210]. Roger [700]. Role [293, 385]. Rolf [260, 232]. Ronald [664, 336, 427]. Rosen [156, 93]. Rosenberg [606, 569]. Rosenblueth [238]. Rosenkrantz [700]. Roy [15]. Royce [286, 309]. Rozeboom [286]. Rudner [467, 718]. Rudolf [262, 229]. Rule [515, 514, 513]. Rules [102, 101, 307]. Runciman [282]. Ruse [372, 222]. Russell [427, 486, 550, 43]. Ryan [211]. Rychlak [34].

S [718, 351, 374, 166, 158, 686, 648, 683, 684, 125, 429, 416]. S-R [416]. Sabat [660]. Safran [307]. Salmon [416, 186]. Saul [139]. Sayre [650, 644].

Sceptical [390]. Scepticism [12]. Scheffler [345, 483, 467, 543]. Scheibe [394]. Schemata [95]. Schlegel [231]. Schlesinger [712, 666]. Schmidt [64]. Schriften [190]. Schütte [64]. Science [26, 273, 71, 127, 560, 34, 96, 231, 89, 510, 31, 118, 609, 561, 58, 518, 280, 260, 185, 390, 345, 438, 228, 30, 378, 125, 713, 483, 716, 124, 189, 188, 282, 397, 235, 684, 715, 238, 285, 590, 32, 283, 351, 424, 107, 228, 334]. Sciences [343, 279, 313, 604, 703, 259, 211, 161]. Scientific [413, 310, 149, 136, 493, 94, 434, 595, 437, 398, 130, 249, 218, 706, 108, 8, 174, 342, 357, 519, 607, 184, 252, 376, 46, 415, 666, 285, 482, 141, 428, 317, 716, 148, 371, 236, 32, 33, 25, 674, 313, 687, 192, 306]. Second [318]. Section [560, 12]. Seeger [125]. Selected [585]. Selection [556, 639, 555, 63]. Selective [358]. Self [97, 98]. Self-Consistency [98]. Sellars [15]. Semantic [96, 477]. Semantics [40]. Seminar [187]. Sense [590]. Sensory [8]. Sentences [79, 387]. Set [403, 105]. Seventeenth [351, 334]. Seventeenth-Century [334]. Seymour [310]. Shackle [374]. Sharp [93]. Shea [607]. Siegfried [289]. Signal [602]. Signed [448]. Significance [697, 118, 451, 583, 355, 581, 685]. Significant [293, 497]. Similar [87]. Simon [339]. Simple [51]. Simple-Minded [51]. Simplicity [599, 481, 482]. Simplified [267]. Simulation [253, 370]. Simulations [369]. Simultaneity [491, 677, 551, 676]. Singular [102]. Sketch [582]. Sklar [520, 674]. Sloman [713]. Slow [643]. Sneed [257]. Snyder [277]. Social [282, 703, 259, 615, 306, 211, 279]. Society [334]. Sociobiology [470]. Sociology [212, 684, 212, 157]. Solipsism [117]. Solution [328, 51]. Solutions [616]. Some [440, 554, 118, 586, 41, 646, 7, 202, 415, 537, 63]. Sommers [123]. Space [671, 139, 58, 425, 696, 533, 674, 506, 335, 520]. Spacetime [657, 520]. Spatial [478]. Special [560, 677, 295, 151, 551, 695, 504, 6, 23]. Species [710, 9]. Specificity [29]. Spector [308]. Spectra [552]. Spicker [685]. Spielman [332]. Sprio [568]. Stanley [502]. State [217, 97, 98]. Statement [95]. Statements [546, 440, 76, 275, 178, 222]. States [635]. Statistical [42, 28, 435, 321, 452, 220, 376, 525, 528, 442, 517, 644, 269, 581]. Statistics [314, 303]. Statistische [589]. Status [292]. Steady [97, 98]. Steady-State [97, 98]. Steel [629]. Stegmuller [466, 589, 208, 82]. Step [328]. Stephen [167, 304, 649, 714]. Steven [81, 236]. Stevenson [223]. Stratigraphical [487]. Straus [165]. Strawson [43]. Strong [595, 253, 370, 369]. Structural [550]. Structure [257, 398, 663, 657, 228, 148]. Structures [714, 59]. Stuart [685, 684]. Studies [125, 716, 106, 229, 717]. Study [426, 714, 605, 459, 143, 307, 378]. Subdoxastic [635]. Subjunctive [494]. Substantial [656]. Sufficiency [278]. Suggestions [500]. Sui [491]. Sum [515, 514, 513]. Supercalifragilistic [80]. Supervenience [614]. Supervenient [597]. Suppe [398]. Suppes [191, 234]. Supposed [279]. Survey [686]. Susan [586]. Swinburne [506]. Sykes [394]. Symbols [54]. Symmetrical [497]. Symmetry [103]. Synchrony [602]. Syntactic [96]. Synthese [37, 66, 84, 171]. System [439]. Systematic [307]. Systems [89, 21, 403, 295, 87, 274, 486, 232]. Szabo [160].

T [312, 605, 373, 396]. Talk [392]. Tarski [617]. Taylor [420]. Technology [26, 560, 563, 334]. Ted [233]. Teleological [1]. Teleology [680, 681, 420, 459, 538, 179]. Temporal [505, 368, 395]. Tension [648]. Term [407]. Terms [256, 197]. Terrell [305]. Testability [435]. Tested [515]. Testing [137, 581]. Tests [451, 355]. Thematic [371]. Theology [35]. Theorem [267]. Theoretical [256, 562, 88, 585, 599, 309, 113, 168, 197, 485]. Theorie [208, 466]. Theoriendynamik [466]. Theorienstrukturen [466]. Theories [270, 256, 413, 333, 616, 511, 386, 493, 434, 358, 435, 595, 387, 4, 437, 404, 130, 672, 228, 482, 669, 2, 675, 148, 500, 60, 143, 25, 40, 591, 398, 209, 141]. Theory [187, 448, 268, 287, 54, 89, 24, 139, 149, 233, 363, 284, 670, 638, 131, 123, 137, 454, 300, 107, 677, 216, 385, 617, 208, 150, 249, 694, 119, 368, 357, 234, 360, 715, 575, 636, 641, 428, 529, 564, 269, 331, 716, 441, 33, 320, 34, 379, 105, 260, 240]. Theory-Change [694]. Theory-Ladenness [300]. Theory/Observation [119]. There [44, 221, 672]. Theses [388]. Thesis [419, 103, 9]. Thiele [64]. Things [576]. Thinking [262, 108]. Thomas [166, 648, 629, 180, 535]. Thomson [665]. Thought [249, 371, 664, 63, 16]. Thoughts [537]. Three [548, 175]. Time [276, 329, 319, 58, 221, 385, 605, 335, 475, 114, 155, 59, 520, 682, 674, 657, 139, 81, 506, 425, 701]. Time-Difference [221]. Time-Structures [59]. Todd [378]. Toraldo [608]. Torretti [678]. Total [27]. Toulmin [304]. Trace [384]. Traces [576]. Tradition [14]. Traditional [106, 342]. Traditions [701]. Transitivity [4]. Translation [356, 437]. Transmitted [28, 637]. Transport [602]. Travel [682]. Travis [206]. Treatment [505]. Trigger [701]. Tristram [685]. Triumph [312]. True [255, 526, 436, 254]. Truth [493, 474, 617, 406, 436, 360]. Tuana [622]. Tuomela [485]. Turing [646]. Turnbull [717]. Two [616, 414, 580, 99, 465, 121, 503, 154, 500]. Types [548, 175, 284].

Uncertainty [686, 247]. Understanding [304]. Unification [309]. Uniformity [35]. Universal [675]. Universality [594]. Universe [426]. Unmasked [271]. Unreasonable [244]. Urning [534]. Urquhart [395]. Use [477, 122, 116]. Uses [448]. Using [28]. Utility [479, 547]. Utz [539].

V [487]. Value [560, 107]. Variables [552]. Velocity [574, 6, 23]. Verbal [240]. Verisimilitude [474]. Vicious [429]. Vicious-Circle [429]. View [148, 254, 25]. Views [174, 674]. Vindication [454]. Vision [203]. Visual [358, 262]. Vita [487]. Vita-Finzi [487]. Volume [67, 145, 241, 325, 381, 443, 507, 570, 651, 720, 304, 584, 285].

W [310, 588, 686, 591, 282, 397, 212, 157, 286, 376, 590, 53]. Waddington [168]. Wahrscheinlichkeit [589]. Wallace [285]. Ward [305]. Was [255].

Watson [236]. Way [574, 6, 23]. Weak [601, 253, 370, 369]. Weber [282]. Well [513]. Well-Confirmed [513]. Westfall [351]. Whewell [33, 255, 254]. Whole [474]. Willard [164, 141]. William [158, 607, 285, 378, 396, 114, 33]. Williams [63]. Wilson [255, 470]. Wissenschaftliche [82]. Wissenschaftstheorie [82, 208]. without [574, 366, 367, 6, 23]. Wolfgang [287, 82, 466, 589]. Wood [15]. Worlds [501]. Wright [420, 275]. Writings [585, 683]. Wrong [148]. Wsevolod [157].

Year [686]. Yehuda [683]. Yolton [53]. Young [662]. Yourgrau [287].

Zupan [422]. Zweiter [466].

References

Ayala:1970:TEE

[1] Francisco J. Ayala. Teleological explanations in evolutionary biology. Philosophy of Science, 37(1):1–15, March 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186024.

Simon:1970:APT

[2] Herbert A. Simon. The axiomatization of physical theories. Philosophy of Science, 37(1):16–26, March 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186025.

Stevenson:1970:II

[3] Charles L. Stevenson. If-Iculties. Philosophy of Science, 37(1):27–49, March 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/186026. See comments

[223].

Hesse:1970:TTC

[4] Mary Hesse. Theories and the transitivity of confirmation. Philosophy of Science, 37(1):50–63, March 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186027.

Stalnaker:1970:PC

[5] Robert C. Stalnaker. Probability and conditionals. Philosophy of Science, 37(1):64–80, March 1970. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186028.

Winnie:1970:SROa

[6] John A. Winnie. Special Relativity without one-way velocity assumptions: Part I. Philosophy of Science, 37(1):81–99, March 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186029.

Lewis:1970:SPP

[7] Douglas Lewis. Some problems of perception. Philosophy of Science, 37 (1):100–113, March 1970. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186030.

Margolis:1970:SRO

[8] Joseph Margolis. Scientific realism, ontology, and the sensory modes. Philosophy of Science, 37(1):114–120, March 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186031.

Munson:1970:BSM

[9] Ronald Munson. Biological species: Mr. Lehman’s thesis. Philosophy of Science, 37(1):121–124, March 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186032.

Lehman:1970:RM

[10] Hugh Lehman. Reply to Munson. Philosophy of Science, 37(1):125–130, March 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/186033.

Lauter:1970:ERL

[11] H. A. Lauter. An examination of Reichenbach on laws. Philosophy of Science, 37(1):131–145, March 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186034.

Sievert:1970:HCS

[12] Don Sievert. Historical cross section laws and Austin’s scepticism in “other minds”. Philosophy of Science, 37(1):146–152, March 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186035.

Bub:1970:BRBa

[13] Jeffrey Bub. Book review: The Philosophy of Quantum Mechanics by D. I. Blokhintsev. Philosophy of Science, 37(1):153–156, March 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186036.

Bub:1970:BRBb

[14] Jeffrey Bub. Book review: Quantum Physics and the Philosophical Tradition by Aage Petersen. Philosophy of Science, 37(1):156–158, March 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186037.

Collins:1970:BRB

[15] James Collins. Book review: Lending a Hand to Hylas by Roy Wood Sellars. Philosophy of Science, 37(1):158–159, March 1970. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186038.

Addis:1970:BRB

[16] Laird Addis. Book review: Thought and Change by Ernest Gellner. Philosophy of Science, 37(1):159–162, March 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186039.

Anonymous:1970:RBa

[17] Anonymous. Recent books. Philosophy of Science, 37(1):163–164, March 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186040.

Anonymous:1970:BMa

[18] Anonymous. Back matter. Philosophy of Science, 37(1):165–182, March 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186041.

Anonymous:1970:FMa

[19] Anonymous. Front matter. Philosophy of Science, 37(1):??, March 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186023.

Miller:1970:FPP

[20] George A. Miller. Four philosophical problems of psycholinguists. Philosophy of Science, 37(2):183–199, June 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186668.

Boden:1970:IPS

[21] Margaret A. Boden. Intentionality and physical systems. Philosophy of Science, 37(2):200–214, June 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186669.

Marley:1970:ACM

[22] A. A. J. Marley. Additive conjoint measurement with respect to a pair of orderings. Philosophy of Science, 37(2):215–222, June 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186670.

Winnie:1970:SROb

[23] John A. Winnie. Special Relativity without one-way velocity assumptions: Part II. Philosophy of Science, 37(2):223–238, June 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186671.

Boland:1970:CET

[24] Lawrence A. Boland. Conventionalism and economic theory. Philosophy of Science, 37(2):239–248, June 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186672.

Wolfson:1970:PVS

[25] Robert J. Wolfson. Points of view, scientific theories and econometric models. Philosophy of Science, 37(2):249–260, June 1970. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186673.

Agassi:1970:PES

[26] Joseph Agassi. Positive evidence in science and technology. Philosophy of Science, 37(2):261–270, June 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186674.

McLaughlin:1970:RTE

[27] Andrew McLaughlin. Rationality and total evidence. Philosophy of Science, 37(2):271–278, June 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186675.

Greeno:1970:ESH

[28] James G. Greeno. Evaluation of statistical hypotheses using information transmitted. Philosophy of Science, 37(2):279–294, June 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186676.

Kyburg:1970:MMS

[29] Henry E. Kyburg, Jr. More on maximal specificity. Philosophy of Science, 37(2):295–300, June 1970. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186677.

Purtill:1970:PS

[30] Richard Purtill. The purpose of science. Philosophy of Science, 37(2):301– 306, June 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186678.

Bushkovitch:1970:PSM

[31] A. V. Bushkovitch. Philosophy of science as a model for all philosophy. Philosophy of Science, 37(2):307–311, June 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186679.

Veatch:1970:BRB

[32] Henry Veatch. Book review: Aspects of Scientific Explanation and Other Essays in the Philosophy of Science by Carl G. Hempel. Philosophy of Science, 37(2):312–314, June 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186680.

Walsh:1970:BRB

[33] Harold T. Walsh. Book review: William Whewell’s Theory of Scientific Method by Robert E. Butts. Philosophy of Science, 37(2):314–315, June 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186681.

Ardila:1970:BRB

[34] Ruben Ardila. Book review: A Philosophy of Science for Personality Theory by Joseph F. Rychlak. Philosophy of Science, 37(2):315–316, June 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186682.

Watson:1970:BRB

[35] Richard A. Watson. Book review: The Principle of Uniformity in Geology, Biology, and Theology: Natural Law and Divine Miracle by R. Hooykaas. Philosophy of Science, 37(2):316–317, June 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186683.

Anonymous:1970:RBb

[36] Anonymous. Recent books. Philosophy of Science, 37(2):317, June 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186684.

Anonymous:1970:ASa

[37] Anonymous. Abstracts from synthese. Philosophy of Science, 37(2):318– 322, June 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186685.

Anonymous:1970:BMb

[38] Anonymous. Back matter. Philosophy of Science, 37(2):??, June 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186686.

Anonymous:1970:FMb

[39] Anonymous. Front matter. Philosophy of Science, 37(2):??, June 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186667.

vanFraassen:1970:EBS

[40] Bas C. van Fraassen. On the extension of Beth’s semantics of physical theories. Philosophy of Science, 37(3):325–339, September 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186462.

Glymour:1970:SPR

[41] Clark Glymour. On some patterns of reduction. Philosophy of Science, 37 (3):340–353, September 1970. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186463.

Giere:1970:OSR

[42] Ronald N. Giere. An orthodox statistical resolution of the paradox of confirmation. Philosophy of Science, 37(3):354–362, September 1970. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186464.

Hochberg:1970:SRK

[43] Herbert Hochberg. Strawson, Russell, and the King of France. Philosophy of Science, 37(3):363–384, September 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186465.

Fisk:1970:TNC

[44] Milton Fisk. Are there necessary connections in nature? Philosophy of Science, 37(3):385–404, September 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186466.

Sicha:1970:CNN

[45] Jeffrey F. Sicha. Counting and the natural numbers. Philosophy of Science, 37(3):405–416, September 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186467.

Omer:1970:DMS

[46] I. A. Omer. On the D-N model of scientific explanation. Philosophy of Science, 37(3):417–433, September 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186468.

Morgan:1970:KDE

[47] Charles G. Morgan. Kim on deductive explanation. Philosophy of Science, 37(3):434–439, September 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186469.

Jackson:1970:RIO

[48] Frank Jackson. A reply to “induction and objectivity”. Philosophy of Science, 37(3):440–443, September 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 186470. See comments [49, 698].

Clendinnen:1970:RJ

[49] F. John Clendinnen. A response to Jackson. Philosophy of Science, 37(3): 444–448, September 1970. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186471.

See [48, 698].

Jackson:1970:RR

[50] Frank Jackson. Reply to a response. Philosophy of Science, 37(3):449–451, September 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186472.

Hullett:1970:SMS

[51] James N. Hullett. On a simple-minded solution. Philosophy of Science, 37 (3):452–454, September 1970. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186473.

Michalos:1970:BRB

[52] Alex C. Michalos. Book review: Robots, Men and Minds by Ludwig von Bertalanffy. Philosophy of Science, 37(3):455, September 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186474.

Wilson:1970:BRB

[53] Fred Wilson. Book review: Metaphysical Analysis by J. W. Yolton. Philosophy of Science, 37(3):455–458, September 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186475.

Beardsley:1970:BRB

[54] Monroe C. Beardsley. Book review: Languages of Art: An Approach to a Theory of Symbols by Nelson Goodman. Philosophy of Science, 37(3): 458–463, September 1970. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186476.

Anonymous:1970:AI

[55] Anonymous. Abstracts from inquiry. Philosophy of Science, 37(3):464– 468, September 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186477.

Anonymous:1970:BMc

[56] Anonymous. Back matter. Philosophy of Science, 37(3):??, September 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186478.

Anonymous:1970:FMc

[57] Anonymous. Front matter. Philosophy of Science, 37(3):??, September 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186461.

Grunbaum:1970:STF

[58] Adolf Grünbaum. Space, time and falsifiability critical exposition and reply to “A Panel Discussion of Grünbaum’s Philosophy of Science”. Philosophy of Science, 37(4):469–588, December 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186137.

Stein:1970:PTS

[59] Howard Stein. On the paradoxical time-structures of Gödel. Philosophy of Science, 37(4):589–601, December 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186138.

Swinburne:1970:CBC

[60] R. G. Swinburne. Choosing between confirmation theories. Philosophy of Science, 37(4):602–613, December 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186139.

Dietz:1970:RHR

[61] Stephens M. Dietz. A remark on Hempel’s replies to his critics. Philosophy of Science, 37(4):614–617, December 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186140.

Mullen:1970:CMP

[62] John D. Mullen. A. C. Michalos’ “postulates of rational preference”. Philosophy of Science, 37(4):618–619, December 1970. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186141.

Wimsatt:1970:BRB

[63] William C. Wimsatt. Book review: Adaptation and Natural Selection: A Critique of Some Current Evolutionary Thought by George C. Williams. Philosophy of Science, 37(4):620–623, December 1970. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186142.

Skidmore:1970:BRB

[64] Arthur Skidmore. Book review: Contributions to Mathematical Logic by H. Arnold Schmidt; Kurt Schütte; Ernst Jochen Thiele. Philosophy of Science, 37(4):623–625, December 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186143.

Anonymous:1970:RBc

[65] Anonymous. Recent books. Philosophy of Science, 37(4):625, December 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186144.

Anonymous:1970:ASb

[66] Anonymous. Abstracts from synthese. Philosophy of Science, 37(4):626– 628, December 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186145.

Anonymous:1970:VI

[67] Anonymous. Volume information. Philosophy of Science, 37(4):629–633, December 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186135.

Anonymous:1970:BMd

[68] Anonymous. Back matter. Philosophy of Science, 37(4):??, December 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186146.

Anonymous:1970:FMd

[69] Anonymous. Front matter. Philosophy of Science, 37(4):??, December 1970. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186136.

Shope:1971:PPE

[70] Robert K. Shope. Physical and psychic energy. Philosophy of Science, 38 (1):1–12, March 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186758.

Alston:1971:PEP

[71] William P. Alston. The place of the explanation of particular facts in science. Philosophy of Science, 38(1):13–34, March 1971. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186759.

Laudan:1971:TRC

[72] Larry Laudan. Towards a reassessment of Comte’s ’Méthode positive’. Philosophy of Science, 38(1):35–53, March 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186760.

Lewis:1971:IIM

[73] David Lewis. Immodest inductive methods. Philosophy of Science, 38(1): 54–63, March 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186761.

Madden:1971:HFF

[74] Edward H. Madden. Hume and the fiery furnace. Philosophy of Science, 38(1):64–78, March 1971. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186762.

Martin:1971:CM

[75] Michael Martin. On the conceivability of mechanism. Philosophy of Science, 38(1):79–86, March 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186763.

Ruse:1971:FSB

[76] Michael E. Ruse. Functional statements in biology. Philosophy of Science, 38(1):87–95, March 1971. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186764.

Girill:1971:FMA

[77] T. R. Girill. Formal models and Achinstein’s “analogies”. Philosophy of Science, 38(1):96–104, March 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186765.

Boland:1971:MEE

[78] Lawrence A. Boland. Methodology as an exercise in economic analysis. Philosophy of Science, 38(1):105–117, March 1971. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186766.

Berg:1971:AAR

[79] Jan Berg. On an argument against reduction sentences. Philosophy of Science, 38(1):118–120, March 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186767.

Scheffler:1971:SRR

[80] Israel Scheffler. Supercalifragilistic reduction: A reply to Jan Berg. Philosophy of Science, 38(1):121, March 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186768.

Fitzgerald:1971:BRB

[81] Paul Fitzgerald. Book review: Fate, Logic, and Time by Steven M. Cahn. Philosophy of Science, 38(1):122–126, March 1971. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186769.

Kockelmans:1971:BRB

[82] Joseph J. Kockelmans. Book review: Probleme und Resultate der Wissenschaftstheorie und Analytische Philosophie. Band I: Wissenschaftliche Erklarung und Begrundung by Wolfgang Stegmüller. Philosophy of Science, 38(1):126–132, March 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186770.

Anonymous:1971:RBa

[83] Anonymous. Recent books. Philosophy of Science, 38(1):133, March 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186771.

Anonymous:1971:AS

[84] Anonymous. Abstracts from synthese. Philosophy of Science, 38(1):134– 137, March 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186772.

Anonymous:1971:BMa

[85] Anonymous. Back matter. Philosophy of Science, 38(1):138–156, March 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186773.

Anonymous:1971:FMa

[86] Anonymous. Front matter. Philosophy of Science, 38(1):??, March 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186757.

Luce:1971:SSD

[87] R. Duncan Luce. Similar systems and dimensionally invariant laws. Philosophy of Science, 38(2):157–169, June 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186775.

Child:1971:TDC

[88] James Child. On the theoretical dependence of correspondence postulates. Philosophy of Science, 38(2):170–177, June 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186776.

Blachowicz:1971:STE

[89] James A. Blachowicz. Systems theory and evolutionary models of the development of science. Philosophy of Science, 38(2):178–199, June 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186777.

Munson:1971:BA

[90] Ronald Munson. Biological adaptation. Philosophy of Science, 38(2):200– 215, June 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186778.

Dretske:1971:RKP

[91] Fred I. Dretske. Reasons, knowledge, and probability. Philosophy of Science, 38(2):216–220, June 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186779.

Lande:1971:DFQ

[92] Alfred Landé. The decline and fall of quantum dualism. Philosophy of Science, 38(2):221–223, June 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186780.

Hooker:1971:SRE

[93] C. A. Hooker. Sharp and the refutation of the Einstein, Podolsky, Rosen Paradox. Philosophy of Science, 38(2):224–233, June 1971. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186781.

Foss:1971:ACB

[94] Laurence Foss. Art as cognitive: Beyond scientific realism. Philosophy of Science, 38(2):234–250, June 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186782.

Nissen:1971:NFS

[95] Lowell Nissen. Neutral functional statement schemata. Philosophy of Science, 38(2):251–257, June 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186783.

Barr:1971:SSA

[96] William F. Barr. A syntactic and semantic analysis of idealizations in science. Philosophy of Science, 38(2):258–272, June 1971. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186784.

Hawkins:1971:SCS

[97] David Hawkins. On the self consistency of the steady-state cosmology. Philosophy of Science, 38(2):273–279, June 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186785.

Schlegel:1971:SCS

[98] Richard Schlegel. On the self-consistency of the steady-state cosmology: Reply to David Hawkins. Philosophy of Science, 38(2):280–281, June 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186786.

Guleserian:1971:TAE

[99] Theodore Guleserian. On two aspects of eliminative materialism. Philosophy of Science, 38(2):282–289, June 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186787.

Butrick:1971:PR

[100] Richard Butrick. Putnam’s revolution. Philosophy of Science, 38(2):290– 292, June 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186788.

Michalos:1971:HRA

[101] Alex C. Michalos. Hilpinen’s rules of acceptance and inductive logic. Philosophy of Science, 38(2):293–302, June 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186789.

Hilpinen:1971:RAI

[102] Risto Hilpinen and Jaakko Hintikka. Rules of acceptance, indices of lawlikeness, and singular inductive inference: Reply to a critical discussion. Philosophy of Science, 38(2):303–307, June 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186790.

Hopson:1971:AST

[103] Ronald C. Hopson. Angel’s symmetry thesis. Philosophy of Science, 38 (2):308–309, June 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186791.

Lucas:1971:MPM

[104] J. R. Lucas. Metamathematics and the philosophy of mind: A rejoinder. Philosophy of Science, 38(2):310–313, June 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186792.

Woods:1971:BRB

[105] John Woods. Book review: Set Theory by K. Kuratowski; A. Mostowski. Philosophy of Science, 38(2):314–315, June 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186793.

Galligan:1971:BRB

[106] Edward M. Galligan and Michael D. Resnik. Book review: Studies on Gottlob Frege and Traditional Philosophy by Ignacio Angelelli. Philosophy of Science, 38(2):316–319, June 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186794.

George:1971:BRB

[107] R. George. Book review: Science and the Theory of Value by Peter Caws. Philosophy of Science, 38(2):319–321, June 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186795.

Madden:1971:BRB

[108] Edward H. Madden. Book review: The Principles of Scientific Thinking by R. Harré. Philosophy of Science, 38(2):321–323, June 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186796.

Anonymous:1971:RBb

[109] Anonymous. Recent books. Philosophy of Science, 38(2):324–325, June 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186797.

Anonymous:1971:BMb

[110] Anonymous. Back matter. Philosophy of Science, 38(2):326, June 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186798.

Anonymous:1971:FMb

[111] Anonymous. Front matter. Philosophy of Science, 38(2):??, June 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186774.

Swinburne:1971:PPE

[112] R. G. Swinburne. The probability of particular events. Philosophy of Science, 38(3):327–343, September 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186006.

Martin:1971:TLP

[113] J. E. Martin. Theoretical languages in psychology. Philosophy of Science, 38(3):344–352, September 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186007.

Myers:1971:WJT

[114] Gerald E. Myers. William James on time perception. Philosophy of Science, 38(3):353–360, September 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186008.

Hanen:1971:CAC

[115] Marsha Hanen. Confirmation and adequacy conditions. Philosophy of Science, 38(3):361–368, September 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186009.

Travis:1971:CUC

[116] Janet L. Travis. A criticism of the use of the concept of “dominant group” in arguments for evolutionary progressivism. Philosophy of Science, 38(3):

369–375, September 1971. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186010.

Mitroff:1971:SEP

[117] Ian I. Mitroff. Solipsism: An essay in psychological philosophy. Philosophy of Science, 38(3):376–394, September 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186011.

Cantore:1971:HSS

[118] Enrico Cantore. Humanistic significance of science: Some methodological considerations. Philosophy of Science, 38(3):395–412, September 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186012.

Machamer:1971:RDT

[119] Peter K. Machamer. A recent drawing of the theory/observation distinction. Philosophy of Science, 38(3):413–414, September 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186013.

Cornman:1971:OWI

[120] James W. Cornman. Observing and what it entails. Philosophy of Science, 38(3):415–417, September 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186014.

Hooker:1971:AKR

[121] C. A. Hooker. Against Krips’ resolution of two paradoxes in quantum mechanics. Philosophy of Science, 38(3):418–428, September 1971. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186015.

Gauthier:1971:UAM

[122] Yvon Gauthier. The use of the axiomatic method in quantum physics. Philosophy of Science, 38(3):429–437, September 1971. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186016.

Englebretsen:1971:STP

[123] George Englebretsen. Sommers’ theory and the paradox of confirmation. Philosophy of Science, 38(3):438–441, September 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186017.

Wilson:1971:ACS

[124] Fred Wilson. On Achinstein’s concepts of science. Philosophy of Science, 38(3):442–452, September 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186018.

Salt:1971:BRB

[125] David Salt. Book review: Ernst Mach, Physicist and Philosopher (Boston Studies in the Philosophy of Science) by Robert S. Cohen; Raymond J. Seeger. Philosophy of Science, 38(3):453–454, September 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186019.

Anonymous:1971:RBc

[126] Anonymous. Recent books. Philosophy of Science, 38(3):454, September 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186020.

Anonymous:1971:ABJ

[127] Anonymous. Abstracts from British journal for the philosophy of science. Philosophy of Science, 38(3):455–462, September 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186021.

Anonymous:1971:BMc

[128] Anonymous. Back matter. Philosophy of Science, 38(3):463–466, September 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/186022.

Anonymous:1971:FMc

[129] Anonymous. Front matter. Philosophy of Science, 38(3):??, September 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186005.

Kordig:1971:CST

[130] Carl R. Kordig. The comparability of scientific theories. Philosophy of Science, 38(4):467–485, December 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186689.

Eberle:1971:ROT

[131] Rolf A. Eberle. Replacing one theory by another under preservation of a given feature. Philosophy of Science, 38(4):486–501, December 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186690.

Hattiangadi:1971:AIC

[132] J. N. Hattiangadi. Alternatives and incommensurables: The case of Darwin and Kelvin. Philosophy of Science, 38(4):502–507, December 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186691.

Gardner:1971:QLR

[133] Michael R. Gardner. Is quantum logic really logic? Philosophy of Science, 38(4):508–529, December 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186692.

Todd:1971:IP

[134] William Todd. Intentions and programs. Philosophy of Science, 38(4): 530–541, December 1971. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186693.

Nickles:1971:CLE

[135] Thomas Nickles. Covering law explanation. Philosophy of Science, 38(4): 542–561, December 1971. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186694.

Carloye:1971:ISM

[136] Jack C. Carloye. An interpretation of scientific models involving analogies. Philosophy of Science, 38(4):562–569, December 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186695.

Firestone:1971:RCF

[137] Joseph M. Firestone. Remarks on concept formation: Theory building and theory testing. Philosophy of Science, 38(4):570–604, December 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186696.

Erlichson:1971:LFC

[138] Herman Erlichson. The Lorentz–Fitzgerald Contraction Hypothesis and the Combined Rod Contraction-Clock Retardation Hypothesis. Philosophy of Science, 38(4):605–609, December 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186697.

Brittan:1971:BRB

[139] Gordon G. Brittan, Jr. Book review: A Deductive Theory of Space and Time by Saul A. Basri. Philosophy of Science, 38(4):610–612, December 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186698.

Michalos:1971:BRB

[140] Alex C. Michalos, Robert E. Butts, and Michael David Resnik. Book review: The Encyclopedia of Philosophy by Paul Edwards. Philosophy of Science, 38(4):612–614, December 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186699.

Ruse:1971:BRB

[141] Michael Ruse. Book review: Anomalies and Scientific Theories by Willard C. Humphreys. Philosophy of Science, 38(4):614–616, December 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186700.

Anonymous:1971:RBd

[142] Anonymous. Recent books. Philosophy of Science, 38(4):616, December 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186702.

Thomas:1971:BRB

[143] Ivo Thomas. Book review: Probability and Opinion. A Study in the Medieval Presuppositions of Post-medieval Theories of Probability by Edmund F. Byrne. Philosophy of Science, 38(4):616, December 1971. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186701.

Anonymous:1971:AI

[144] Anonymous. Abstracts from inquiry. Philosophy of Science, 38(4):617– 624, December 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186703.

Anonymous:1971:VI

[145] Anonymous. Volume information. Philosophy of Science, 38(4):625–629, December 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186687.

Anonymous:1971:BMd

[146] Anonymous. Back matter. Philosophy of Science, 38(4):??, December 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186704.

Anonymous:1971:FMd

[147] Anonymous. Front matter. Philosophy of Science, 38(4):??, December 1971. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186688.

Suppe:1972:WWR

[148] Frederick Suppe. “What’s wrong with the received view on the structure of scientific theories?”. Philosophy of Science, 39(1):1–19, March 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186589.

Brody:1972:TAT

[149] B. A. Brody. Towards an Aristotelean theory of scientific explanation. Philosophy of Science, 39(1):20–31, March 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186590.

Krause:1972:AFC

[150] Merton S. Krause. An analysis of Festinger’s cognitive dissonance theory. Philosophy of Science, 39(1):32–50, March 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186591.

Gutting:1972:EDS

[151] Gary Gutting. Einstein’s discovery of Special Relativity. Philosophy of Science, 39(1):51–68, March 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186592.

Dretske:1972:CI

[152] Fred I. Dretske and Aaron Snyder. Causal irregularity. Philosophy of Science, 39(1):69–71, March 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186593.

Savitt:1972:CC

[153] Steven Savitt. I-counting is counting. Philosophy of Science, 39(1):72– 73, March 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186594.

Morgan:1972:TPM

[154] Charles G. Morgan. On two proposed models of explanation. Philosophy of Science, 39(1):74–81, March 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186595.

Ruse:1972:BT

[155] Michael Ruse. Bunge on time. Philosophy of Science, 39(1):82, March 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186596.

Erlichson:1972:EPR

[156] Herman Erlichson. The Einstein–Podolski–Rosen Paradox. Philosophy of Science, 39(1):83–85, March 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186597.

Michalos:1972:BRBa

[157] Alex C. Michalos. Book review: Causation and Functionalism in Sociology by Wsevolod W. Isajiw. Philosophy of Science, 39(1):86–88, March 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186598.

Corcoran:1972:BRBa

[158] John Corcoran. Book review: Foundations of Mathematics by William S. Hatcher. Philosophy of Science, 39(1):88–90, March 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186599.

Michalos:1972:BRBb

[159] Alex C. Michalos. Book review: The Problem of Inductive Logic by Imre Lakatos. Philosophy of Science, 39(1):90–91, March 1972. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186600.

vanHeijenoort:1972:BRB

[160] J. van Heijenoort. Book review: The Collected Papers of Gerhard Gentzen by M. E. Szabo. Philosophy of Science, 39(1):91, March 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186601.

Ruse:1972:BRBa

[161] Michael E. Ruse. Book review: The Relations between the Sciences by C. F. A. Pantin. Philosophy of Science, 39(1):91–92, March 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186602.

Lehman:1972:BRBa

[162] Hugh Lehman. Book review: Criticism and the Growth of Knowledge by Imre Lakatos; Alan Musgrave. Philosophy of Science, 39(1):92–95, March 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186603.

Chabdack:1972:BRB

[163] T. C. Chabdack. Book review: Psychological Explanation by Jerry A. Fodor. Philosophy of Science, 39(1):95–97, March 1972. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186604.

Corcoran:1972:BRBb

[164] John Corcoran. Book review: Philosophy of Logic by Willard Van Orman Quine. Philosophy of Science, 39(1):97–99, March 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186605.

Seligman:1972:BRB

[165] Paul Seligman. Book review: Psychiatry and Philosophy by Erwin Straus; Maurice Natanson; Henri Ey. Philosophy of Science, 39(1):99–101, March 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186606.

Coleman:1972:BRB

[166] William Coleman. Book review: Ideas of Life and Matter by Thomas S. Hall. Philosophy of Science, 39(1):101–102, March 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186607.

Brody:1972:BRB

[167] Boruch A. Brody. Book review: The Legacy of Logical Positivism by Peter Achenstein; Stephen Barker. Philosophy of Science, 39(1):102–104, March 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186608.

Ruse:1972:BRBb

[168] Michael Ruse. Book review: Towards a Theoretical Biology by C. H. Waddington. Philosophy of Science, 39(1):105–106, March 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186609.

Corcoran:1972:BRBc

[169] John Corcoran. Book review: Hilbert by Constance Reid. Philosophy of Science, 39(1):106–108, March 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186610.

Anonymous:1972:RBa

[170] Anonymous. Recent books. Philosophy of Science, 39(1):109, March 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186611.

Anonymous:1972:AS

[171] Anonymous. Abstracts from synthese. Philosophy of Science, 39(1):110– 112, March 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186612.

Anonymous:1972:BMa

[172] Anonymous. Back matter. Philosophy of Science, 39(1):113–130, March 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186613.

Anonymous:1972:FMa

[173] Anonymous. Front matter. Philosophy of Science, 39(1):??, March 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186588.

Maxwell:1972:CPV

[174] Nicholas Maxwell. A critique of Popper’s views on scientific method. Philosophy of Science, 39(2):131–152, June 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186717.

Sharvy:1972:TTR

[175] Richard Sharvy. Three types of referential opacity. Philosophy of Science, 39(2):153–161, June 1972. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186718.

Uchii:1972:ILC

[176] Soshichi Uchii. Inductive logic with causal modalities: A probabilistic approach. Philosophy of Science, 39(2):162–178, June 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186719.

Lincicome:1972:RED

[177] David Lincicome. Reporting, evaluating, describing. Philosophy of Science, 39(2):179–194, June 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186720.

Scoledes:1972:DQM

[178] Aristotle G. M. Scoledes. The determinism of quantum-mechanical probability statements. Philosophy of Science, 39(2):195–203, June 1972. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186721.

Wright:1972:ET

[179] Larry Wright. Explanation and teleology. Philosophy of Science, 39(2): 204–218, June 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186722.

Daniels:1972:TRD

[180] Norman Daniels. Thomas Reid’s discovery of a non-Euclidean geometry. Philosophy of Science, 39(2):219–234, June 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186723.

Achinstein:1972:MAR

[181] Peter Achinstein. Models and analogies: A reply to Girill. Philosophy of Science, 39(2):235–240, June 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186724.

Girill:1972:AMR

[182] T. R. Girill. Analogies and models revisited. Philosophy of Science, 39(2): 241–244, June 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186725.

Schwartz:1972:PP

[183] Robert Schwartz. Paradox and projection. Philosophy of Science, 39(2): 245–248, June 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186726.

Miettinen:1972:OMS

[184] Seppo K. Miettinen. On Omer’s model of scientific explanation. Philosophy of Science, 39(2):249–251, June 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186727.

Laszlo:1972:EIC

[185] E. Laszlo and H. Margenau. The emergence of integrative concepts in contemporary science. Philosophy of Science, 39(2):252–259, June 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186728.

Torretti:1972:RSP

[186] Roberto Torretti. Remarks on Salmon’s Paradox of Primes. Philosophy of Science, 39(2):260–262, June 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186729.

Agassi:1972:BRI

[187] Joseph Agassi. Book review: The interface of philosophy and physics: Delaware Seminar in the Foundations of Physics, and Quantum Theory and Reality. Philosophy of Science, 39(2):263–265, June 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186730.

Bunge:1972:BRB

[188] Mario Bunge. Book review: Mach’s Philosophy of Science by J. Bradley. Philosophy of Science, 39(2):266–267, June 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186731.

Bradie:1972:BRB

[189] Michael P. Bradie and James D. Stuart. Book review: Metaphysics and the Philosophy of Science by Gerd Buchdahl. Philosophy of Science, 39 (2):267–269, June 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186732.

Resnik:1972:BRBa

[190] Michael D. Resnik. Book review: Nachgelassene Schriften by Gottlob Frege; Hans Hermes; Friedrich Kambartel; Friedrich Kaulbach. Philosophy of Science, 39(2):269–271, June 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186733.

Michalos:1972:BRBc

[191] Alex C. Michalos. Book review: Information and Inference by J. Hintikka; P. Suppes. Philosophy of Science, 39(2):271–272, June 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186734.

Lehman:1972:BRBb

[192] Hugh Lehman. Book review: Scientific Explanation by Nicholas Rescher. Philosophy of Science, 39(2):272–274, June 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186735.

Anonymous:1972:RBb

[193] Anonymous. Recent books. Philosophy of Science, 39(2):275–276, June 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186736.

Anonymous:1972:AI

[194] Anonymous. Abstracts from inquiry. Philosophy of Science, 39(2):277– 283, June 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186737.

Anonymous:1972:BMb

[195] Anonymous. Back matter. Philosophy of Science, 39(2):284, June 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186738.

Anonymous:1972:FMb

[196] Anonymous. Front matter. Philosophy of Science, 39(2):??, June 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186716.

Sinks:1972:FET

[197] John D. Sinks. Fictionalism and the elimination of theoretical terms. Philosophy of Science, 39(3):285–290, September 1972. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186148.

Globus:1972:BFP

[198] Gordon G. Globus. Biological foundations of the psychoneural identity hypothesis. Philosophy of Science, 39(3):291–301, September 1972. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186149.

Pollock:1972:LP

[199] John L. Pollock. The logic of projectibility. Philosophy of Science, 39(3): 302–314, September 1972. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186150.

Phillips:1972:DI

[200] John N. Phillips. Degrees of interpretation. Philosophy of Science, 39(3): 315–321, September 1972. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186151.

Glymour:1972:PC

[201] Clark Glymour. Physics by convention. Philosophy of Science, 39(3): 322–340, September 1972. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186152.

Macklin:1972:MHM

[202] Ruth Macklin. Mental health and mental illness: Some problems of definition and concept formation. Philosophy of Science, 39(3):341–365, September 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186153.

Scheffler:1972:VRP

[203] Israel Scheffler. Vision and revolution: A postscript on Kuhn. Philosophy of Science, 39(3):366–374, September 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186154.

Spielman:1972:LII

[204] Stephen Spielman. Lewis on immodest inductive models. Philosophy of Science, 39(3):375–377, September 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186155.

Varela:1972:MBE

[205] Francisco Varela and Humberto Maturana. Mechanism and biological explanation. Philosophy of Science, 39(3):378–382, September 1972. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186156.

Glickman:1972:BPD

[206] Maurice J. A. Glickman. Biological progress and dominance: A reply to Janet L. Travis. Philosophy of Science, 39(3):383–387, September 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186157.

Gram:1972:CDR

[207] M. S. Gram. Causation and direct realism. Philosophy of Science, 39(3): 388–396, September 1972. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186158.

Kockelmans:1972:DRS

[208] Joseph J. Kockelmans. Discussion review: Stegmüller on the relationship between theory and experience: Theorie und Erfahrung: Probleme und Resultate der Wissenschaftstheorie und Analytische Philosophie. Philosophy of Science, 39(3):397–420, September 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186159.

Resnik:1972:BRBb

[209] Michael David Resnik. Book review: The Logic of Empirical Theories by Marian Przelecki. Philosophy of Science, 39(3):421–423, September 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186160.

Boland:1972:BRB

[210] L. A. Boland. Book review: The Entropy Law and the Economic Process by Nicholas Georgescu-Roegen. Philosophy of Science, 39(3):423–424, September 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186161.

Rosenberg:1972:BRBa

[211] Alexander Rosenberg. Book review: The Philosophy of the Social Sciences by Alan Ryan. Philosophy of Science, 39(3):424–427, September 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186162.

McLaughlin:1972:BRB

[212] Andrew McLaughlin. Book review: A Sociology of Sociology by Robert W. Friedrichs. Philosophy of Science, 39(3):427–429, September 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186163.

Anonymous:1972:BMc

[213] Anonymous. Back matter. Philosophy of Science, 39(3):430, September 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186164.

Anonymous:1972:FMc

[214] Anonymous. Front matter. Philosophy of Science, 39(3):??, September 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186147.

Boyd:1972:DLP

[215] Richard Boyd. Determinism, laws, and predictability in principle. Philosophy of Science, 39(4):431–450, December 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186328.

Grossman:1972:QMI

[216] Neal Grossman. Quantum mechanics and interpretations of probability theory. Philosophy of Science, 39(4):451–460, December 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186329.

Daly:1972:AAE

[217] Richard Daly. On arguments against the empirical adequacy of finite state grammar. Philosophy of Science, 39(4):461–475, December 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186330.

Leplin:1972:CFS

[218] Jarrett Leplin. Contextual falsification and scientific methodology. Philosophy of Science, 39(4):476–490, December 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186331.

Hull:1972:RGB

[219] David L. Hull. Reduction in genetics — biology or philosophy? Philosophy of Science, 39(4):491–499, December 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186332.

Lehman:1972:SE

[220] Hugh Lehman. Statistical explanation. Philosophy of Science, 39(4):500– 506, December 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186333.

Hedman:1972:WTM

[221] Carl G. Hedman. On when there must be a time-difference between cause and effect. Philosophy of Science, 39(4):507–511, December 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186334.

Wright:1972:CRA

[222] Larry Wright. A comment on Ruse’s analysis of function statements. Philosophy of Science, 39(4):512–514, December 1972. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186335.

Martin:1972:SII

[223] R. M. Martin. On Stevenson’s If-Iculties. Philosophy of Science, 39(4): 515–521, December 1972. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186336.

See [3].

Berent:1972:DPI

[224] Paul Berent. Disconfirmation by positive instances. Philosophy of Science, 39(4):522, December 1972. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186337.

Rose:1972:CCI

[225] Lynn E. Rose. Countering a counter-intuitive probability. Philosophy of Science, 39(4):523–524, December 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186338.

Ruse:1972:BA

[226] Michael Ruse. Biological adaptation. Philosophy of Science, 39(4):525– 528, December 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186339.

Munson:1972:BAR

[227] Ronald Munson. Biological adaptation: A reply. Philosophy of Science, 39 (4):529–532, December 1972. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186340.

Nowak:1972:LST

[228] Leszek Nowak. Laws of science, theories, measurement: (comments on Ernest Nagel’s The Structure of Science). Philosophy of Science, 39(4): 533–548, December 1972. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186341.

Kyburg:1972:BRB

[229] Henry E. Kyburg, Jr. Book review: Studies in Inductive Logic and Probability by Rudolf Carnap; Richard C. Jeffrey. Philosophy of Science, 39 (4):549–553, December 1972. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186342.

Adams:1972:BRB

[230] Ernest W. Adams. Book review: Geometry and Chronometry in Philosophical Perspective by Adolf Grünbaum. Philosophy of Science, 39(4): 553–555, December 1972. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186343.

Barr:1972:BRB

[231] William F. Barr. Book review: Inquiry into Science: Its Domain and Limits by Richard Schlegel. Philosophy of Science, 39(4):555–556, December 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186344.

Wilson:1972:BRB

[232] Fred Wilson. Book review: Nominalistic Systems by Rolf A. Eberle. Philosophy of Science, 39(4):556–558, December 1972. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186345.

Cartwright:1972:BRB

[233] Nancy Cartwright. Book review: Quantum Theory and beyond by Ted Bastin. Philosophy of Science, 39(4):558–560, December 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186346.

Michalos:1972:BRBd

[234] Alex C. Michalos. Book review: A Probabilistic Theory of Causality by P. Suppes. Philosophy of Science, 39(4):560–561, December 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186347.

Massey:1972:BRB

[235] Gerald J. Massey. Book review: An Introduction to the Philosophy of Science by Karel Lambert; Gordon G. Brittan. Philosophy of Science, 39 (4):561–564, December 1972. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186348.

Tuggle:1972:BRB

[236] H. David Tuggle. Book review: Explanation in Archeology; An Explicitly Scientific Approach by Patty Jo Watson; Steven A. Leblanc; Charles L. Redman. Philosophy of Science, 39(4):564–566, December 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186349.

Michalos:1972:BRBe

[237] Alex C. Michalos. Book review: The Implications of Induction by L.J. Cohen. Philosophy of Science, 39(4):566–567, December 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186350.

Puccetti:1972:BRB

[238] Roland Puccetti. Book review: Mind and Brain: A Philosophy of Science by Arturo Rosenblueth. Philosophy of Science, 39(4):567–569, December 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186351.

Martin:1972:BRB

[239] Michael Martin. Book review: Fact and Relevance: Essays on Historical Method by M. M. Postan. Philosophy of Science, 39(4):569–572, December 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186352.

Rosenberg:1972:BRBb

[240] Alexander Rosenberg. Book review: Theory Construction: From Verbal to Mathematical Formulations by Hubert M. Blalock. Philosophy of Science, 39(4):572–573, December 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186353.

Anonymous:1972:VI

[241] Anonymous. Volume information. Philosophy of Science, 39(4):575–580, December 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186326.

Anonymous:1972:BMd

[242] Anonymous. Back matter. Philosophy of Science, 39(4):??, December 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186354.

Anonymous:1972:FMd

[243] Anonymous. Front matter. Philosophy of Science, 39(4):??, December 1972. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186327.

Schagrin:1973:BU

[244] Morton L. Schagrin. On being unreasonable. Philosophy of Science, 40(1): 1–9, March 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186356.

Byerly:1973:RFM

[245] Henry C. Byerly and Vincent A. Lazara. Realist foundations of measurement. Philosophy of Science, 40(1):10–28, March 1973. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186357.

Kochanski:1973:CLP

[246] Zdzislaw Kochanski. Conditions and limitations of prediction-making in biology. Philosophy of Science, 40(1):29–51, March 1973. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186358.

Kirschenmann:1973:RUR

[247] Peter Kirschenmann. Reciprocity in the uncertainty relations. Philosophy of Science, 40(1):52–58, March 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186359.

Rozeboom:1973:DR

[248] William W. Rozeboom. Dispositions revisited. Philosophy of Science, 40(1):59–74, March 1973. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186360.

Laszlo:1973:IST

[249] Ervin Laszlo. The ideal scientific theory: A thought experiment. Philosophy of Science, 40(1):75–87, March 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186361.

Brownstein:1973:BP

[250] Donald Brownstein. Basic particulars. Philosophy of Science, 40(1):88– 96, March 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186362.

Romanos:1973:RP

[251] George D. Romanos. Reflexive predictions. Philosophy of Science, 40(1): 97–109, March 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186363.

Morgan:1973:OSE

[252] Charles G. Morgan. Omer on scientific explanation. Philosophy of Science, 40(1):110–117, March 1973. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186364.

Rosenberg:1973:FDB

[253] A. Rosenberg and N. J. Mackintosh. On Fodor’s distinction between strong and weak equivalence in machine simulation. Philosophy of Science, 40(1):118–120, March 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186365.

Wilson:1973:BWV

[254] David B. Wilson. Butts on Whewell’s view of true causes. Philosophy of Science, 40(1):121–124, March 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186366.

Butts:1973:RDW

[255] Robert E. Butts. Reply to David Wilson: Was Whewell interested in true causes? Philosophy of Science, 40(1):125–128, March 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186367.

Berent:1973:TTI

[256] Paul Berent. Theoretical terms in infinite theories. Philosophy of Science, 40(1):129, March 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186368.

Hooker:1973:BRB

[257] C. A. Hooker. Book review: The Logical Structure of Mathematical Physics by Joseph D. Sneed. Philosophy of Science, 40(1):130–131, March 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186369.

Corcoran:1973:BRBa

[258] John Corcoran. Book review: Philosophy of Logic by Hilary Putnam. Philosophy of Science, 40(1):131–133, March 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186370.

McLaughlin:1973:BRB

[259] Andrew McLaughlin. Book review: Patterns of Discovery in the Social Sciences by Paul Diesing. Philosophy of Science, 40(1):133–136, March 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186371.

Kockelmans:1973:BRB

[260] Joseph J. Kockelmans. Book review: Theory of Science. Attempt at a Detailed and in the Main Novel Exposition of Logic with Constant Attention to Earlier Authors by Bernard Bolzano; Rolf George. Philosophy of Science, 40(1):136–141, March 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186372.

Michalos:1973:BRB

[261] Alex C. Michalos. Book review: The Matter of Chance by D. H. Mellor. Philosophy of Science, 40(1):141, March 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186373.

Bornstein:1973:BRB

[262] Marc Bornstein. Book review: Visual Thinking by Rudolf Arnheim. Philosophy of Science, 40(1):141–144, March 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186374.

Anonymous:1973:RBa

[263] Anonymous. Recent books. Philosophy of Science, 40(1):145–147, March 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186375.

Anonymous:1973:AI

[264] Anonymous. Abstracts from inquiry. Philosophy of Science, 40(1):148– 152, March 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186376.

Anonymous:1973:BMa

[265] Anonymous. Back matter. Philosophy of Science, 40(1):153–173, March 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186377.

Anonymous:1973:FMa

[266] Anonymous. Front matter. Philosophy of Science, 40(1):??, March 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186355.

Mackay:1973:SPI

[267] Alfred F. Mackay. A simplified proof of an impossibility theorem. Philosophy of Science, 40(2):175–177, June 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186379.

Armstrong:1973:EFM

[268] D. M. Armstrong. Epistemological foundations for a materialist theory of the mind. Philosophy of Science, 40(2):178–193, June 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186380.

Sklar:1973:SEE

[269] Lawrence Sklar. Statistical explanation and ergodic theory. Philosophy of Science, 40(2):194–212, June 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186381.

Alder:1973:T

[270] Michael D. Alder. On theories. Philosophy of Science, 40(2):213–226, June 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186382.

Rosenkrantz:1973:PMU

[271] R. D. Rosenkrantz. Probability magic unmasked. Philosophy of Science, 40(2):227–233, June 1973. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186383.

Noren:1973:PME

[272] Stephen J. Noren. The picturability of micro-entities. Philosophy of Science, 40(2):234–241, June 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186384.

Almeder:1973:SI

[273] Robert F. Almeder. Science and idealism. Philosophy of Science, 40(2): 242–254, June 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186385.

Mitroff:1973:SIM

[274] Ian I. Mitroff. Systems, inquiry, and the meanings of falsification. Philosophy of Science, 40(2):255–276, June 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186386.

Ruse:1973:RWA

[275] Michael Ruse. A reply to Wright’s analysis of functional statements. Philosophy of Science, 40(2):277–280, June 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186387.

Clatterbaugh:1973:NNT

[276] Kenneth C. Clatterbaugh. A note on Newtonian time. Philosophy of Science, 40(2):281–284, June 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186388.

Beauchamp:1973:CIR

[277] Tom L. Beauchamp. On causal irregularity: A reply to Dretske and Snyder. Philosophy of Science, 40(2):285–287, June 1973. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186389.

Dretske:1973:CSR

[278] Fred I. Dretske and Aaron Snyder. Causality and sufficiency: Reply to Beauchamp. Philosophy of Science, 40(2):288–291, June 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186390.

Vetterling:1973:SMD

[279] Mary K. Vetterling. On a supposed methodological difference between the natural and social sciences. Philosophy of Science, 40(2):292–293, June 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186391.

Hoy:1973:CAA

[280] Ronald C. Hoy. Cognitive aspects of art and science. Philosophy of Science, 40(2):294–297, June 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186392.

Foss:1973:DDJ

[281] Laurence Foss. Does Don Juan really fly? Philosophy of Science, 40(2): 298–316, June 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186393.

Heil:1973:BRB

[282] John Heil. Book review: A Critique of Max Weber’s Philosophy of Social Science by W. G. Runciman. Philosophy of Science, 40(2):317–318, June 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186394.

Armstrong:1973:BRB

[283] H. L. Armstrong. Book review: Science at the Crossroads by Herbert Dingle. Philosophy of Science, 40(2):318–319, June 1973. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186395.

Corcoran:1973:BRBb

[284] John Corcoran and John Richards. Book review: The Theory of Logical Types by Irving M. Copi. Philosophy of Science, 40(2):319–321, June 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186396.

Rice:1973:BRB

[285] Lee C. Rice. Book review: Causality and Scientific Explanation. Volume I. Medieval and Early Classical Science by William A. Wallace. Philosophy of Science, 40(2):321–322, June 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186397.

Moore:1973:BRB

[286] H. F. Moore. Book review: The Psychology of Knowing by J. R. Royce; W. W. Rozeboom. Philosophy of Science, 40(2):322–323, June 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186398.

Audi:1973:BRB

[287] Michael N. Audi. Book review: Perspectives in Quantum Theory: Essays in Honor of Alfred Landé by Wolfgang Yourgrau; Alwyn Van Der Merwe. Philosophy of Science, 40(2):323–324, June 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186399.

MacCormac:1973:BRB

[288] Earl R. MacCormac. Book review: The Concept of Meaninglessness by Edward Erwin. Philosophy of Science, 40(2):324–326, June 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186400.

Margenau:1973:BRB

[289] Henry Margenau. Book review: Protophysik by Siegfried Müller-Markus. Philosophy of Science, 40(2):326–327, June 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186401.

Anonymous:1973:BMb

[290] Anonymous. Back matter. Philosophy of Science, 40(2):??, June 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186402.

Anonymous:1973:FMb

[291] Anonymous. Front matter. Philosophy of Science, 40(2):??, June 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186378.

Earman:1973:MSN

[292] J. Earman and M. Friedman. The meaning and status of Newton’s law of inertia and the nature of gravitational forces. Philosophy of Science, 40 (3):329–359, September 1973. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186195.

Dorling:1973:DIS

[293] Jon Dorling. Demonstrative induction: Its significant role in the history of physics. Philosophy of Science, 40(3):360–372, September 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186196.

Uchii:1973:HOP

[294] Soshichi Uchii. Higher order probabilities and coherence. Philosophy of Science, 40(3):373–381, September 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186197.

Giannoni:1973:SRA

[295] Carlo B. Giannoni. Special Relativity in accelerated systems. Philosophy of Science, 40(3):382–392, September 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186198.

Stemmer:1973:BDE

[296] Nathan Stemmer. Brody’s defense of essentialism. Philosophy of Science, 40(3):393–396, September 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186199.

Salmon:1973:RL

[297] Wesley C. Salmon. Reply to Lehman. Philosophy of Science, 40(3):397– 402, September 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186200.

Steinberg:1973:NIC

[298] Danny Steinberg. Nickles on intensionality and the covering law model. Philosophy of Science, 40(3):403–407, September 1973. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186201.

Nickles:1973:EDR

[299] Thomas Nickles. Explanation and description-relativity. Philosophy of Science, 40(3):408–414, September 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186202.

Gale:1973:KTL

[300] George Gale and Edward Walter. Kordig and the theory-ladenness of observation. Philosophy of Science, 40(3):415–432, September 1973. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186203.

Doppelt:1973:DCP

[301] Gerald Doppelt. Dretske’s conception of perception and knowledge. Philosophy of Science, 40(3):433–446, September 1973. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186204.

Martin:1973:OM

[302] Michael Martin. The objectivity of a methodology. Philosophy of Science, 40(3):447–450, September 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186205.

Kyburg:1973:BRB

[303] Henry E. Kyburg, Jr. Book review: Probability, Induction and Statistics by Bruno de Finetti. Philosophy of Science, 40(3):451–453, September 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186206.

Haworth:1973:BRB

[304] L. L. Haworth. Book review: Human Understanding. Volume I: General Introduction and Part I by Stephen Toulmin. Philosophy of Science, 40 (3):453–454, September 1973. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186207.

Corcoran:1973:BRBc

[305] John Corcoran and David Levin. Book review: Conceptual Notation and Related Articles by Gottlob Frege; Terrell Ward Bynum. Philosophy of Science, 40(3):454–455, September 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186208.

Moor:1973:BRB

[306] James H. Moor. Book review: Scientific Knowledge and Its Social Problems by Jerome R. Ravetz. Philosophy of Science, 40(3):455–457, September 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/186209.

Brett:1973:BRB

[307] Nathan Brett. Book review: Rules: A Systematic Study by Joan Safran Ganz. Philosophy of Science, 40(3):457–459, September 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186210.

Kirschenmann:1973:BRB

[308] Peter Kirschenmann. Book review: Methodological Foundations of Relativistic Mechanics by Marshall Spector. Philosophy of Science, 40(3): 459–461, September 1973. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186211.

Kitchener:1973:BRB

[309] Richard F. Kitchener. Book review: Toward Unification in Psychology: The First Banff Conference on Theoretical Psychology by Joseph R. Royce. Philosophy of Science, 40(3):461–463, September 1973. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186212.

Blarer:1973:BRB

[310] Joseph Blarer. Book review: Ernst Cassirer: Scientific Knowledge and the Concept of Man by Seymour W. Itzkoff. Philosophy of Science, 40(3): 463–464, September 1973. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186213.

Metzger:1973:BRB

[311] Robert S. Metzger. Book review: Problems in the Foundations of Physics by Mario Bunge. Philosophy of Science, 40(3):464–466, September 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186214.

Blum:1973:BRB

[312] Gary Blum. Book review: The Triumph of the Darwinian Method by Michael T. Ghiselin. Philosophy of Science, 40(3):466–467, September 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186215.

Bradie:1973:BRB

[313] Michael Bradie. Book review: Scientific Method: The Hypothetico-Experimental Laboratory Procedure of the Physical Sciences by James K. Feibleman. Philosophy of Science, 40(3):467–468, September 1973. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186216.

Brier:1973:BRB

[314] Bob Brier. Book review: Randomness, Statistics and Emergence by Philip McShane. Philosophy of Science, 40(3):468–469, September 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186217.

Anonymous:1973:BMc

[315] Anonymous. Back matter. Philosophy of Science, 40(3):??, September 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186218.

Anonymous:1973:FMc

[316] Anonymous. Front matter. Philosophy of Science, 40(3):??, September 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186194.

Simon:1973:DSD

[317] Herbert A. Simon. Does scientific discovery have a logic? Philosophy of Science, 40(4):471–480, December 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186282.

Krantz:1973:FMF

[318] David H. Krantz. Fundamental measurement of force and Newton’s first and second laws of motion. Philosophy of Science, 40(4):481–495, December 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186283.

Freundlich:1973:BAT

[319] Yehudah Freundlich. “becoming” and the asymmetries of time. Philosophy of Science, 40(4):496–517, December 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186284.

Williams:1973:FPE

[320] Mary B. Williams. Falsifiable predictions of evolutionary theory. Philosophy of Science, 40(4):518–537, December 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186285.

Keuth:1973:PPR

[321] Herbert Keuth. On prior probabilities of rejecting statistical hypotheses. Philosophy of Science, 40(4):538–546, December 1973. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186286.

Campbell:1973:APC

[322] Richmond Campbell and Alexander Rosenberg. Action, purpose, and consciousness among the computers. Philosophy of Science, 40(4):547– 557, December 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186287.

Kordig:1973:OI

[323] Carl R. Kordig. Observational invariance. Philosophy of Science, 40(4): 558–569, December 1973. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186288.

Anonymous:1973:RBb

[324] Anonymous. Recent books. Philosophy of Science, 40(4):569, December 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186289.

Anonymous:1973:VI

[325] Anonymous. Volume information. Philosophy of Science, 40(4):571–574, December 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186280.

Anonymous:1973:BMd

[326] Anonymous. Back matter. Philosophy of Science, 40(4):??, December 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186290.

Anonymous:1973:FMd

[327] Anonymous. Front matter. Philosophy of Science, 40(4):??, December 1973. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186281.

Bohnert:1974:LLM

[328] Herbert G. Bohnert. The logico-linguistic mind-brain problem and a proposed step towards its solution. Philosophy of Science, 41(1):1–14, March 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187374.

Earman:1974:AAL

[329] John Earman. An attempt to add a little direction to “The Problem of the Direction of Time”. Philosophy of Science, 41(1):15–47, March 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187375.

Barr:1974:PAI

[330] William F. Barr. A pragmatic analysis of idealizations in physics. Philosophy of Science, 41(1):48–64, March 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187376.

Stenner:1974:TTE

[331] Alfred J. Stenner. Toward a theory of event identity. Philosophy of Science, 41(1):65–83, March 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187377.

Lewis:1974:SLI

[332] David Lewis. Spielman and Lewis on inductive immodesty. Philosophy of Science, 41(1):84–85, March 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187378.

Bushkovitch:1974:MTK

[333] A. V. Bushkovitch. Models, theories, and Kant. Philosophy of Science, 41(1):86–88, March 1974. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187379.

Shea:1974:BRB

[334] William R. Shea. Book review: Science, Technology and Society in Seventeenth-Century England by Robert K. Merton. Philosophy of Science, 41(1):89–90, March 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187380.

Massey:1974:BRB

[335] Gerald J. Massey. Book review: An Introduction to the Philosophy of Time and Space by Bas C. van Fraassen. Philosophy of Science, 41(1):90– 92, March 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187381.

Humphries:1974:BRB

[336] Jill Humphries. Book review: Man and Nature, Philosophical Issues in Biology by Ronald Munson. Philosophy of Science, 41(1):92–94, March 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187382.

Ashworth:1974:BRBa

[337] E. J. Ashworth. Book review: Leibniz’s Philosophy of Logic and Language by Hide Ishiguro. Philosophy of Science, 41(1):94–96, March 1974. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187383.

Moreland:1974:BRB

[338] John M. Moreland. Book review: Mental Images: A Defence by Alastair Hannay. Philosophy of Science, 41(1):96–98, March 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187384.

Morgan:1974:BRBa

[339] Charles G. Morgan. Book review: Reason and Prediction by Simon Blackburn. Philosophy of Science, 41(1):98–100, March 1974. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187385.

Anonymous:1974:BMa

[340] Anonymous. Back matter. Philosophy of Science, 41(1):101–122, March 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187386.

Anonymous:1974:FMa

[341] Anonymous. Front matter. Philosophy of Science, 41(1):??, March 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187373.

Maxwell:1974:RSDa

[342] Nicholas Maxwell. The rationality of scientific discovery Part I: The traditional rationality problem. Philosophy of Science, 41(2):123–153, June 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186864.

Quay:1974:PDC

[343] Paul M. Quay. Progress as a demarcation criterion for the sciences. Philosophy of Science, 41(2):154–170, June 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186865.

Laing:1974:MDC

[344] Richard Laing. Maxwell’s demon and computation. Philosophy of Science, 41(2):171–178, June 1974. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186866.

Meiland:1974:KSO

[345] Jack W. Meiland. Kuhn, Scheffler, and objectivity in science. Philosophy of Science, 41(2):179–187, June 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186867.

Thorpe:1974:QMD

[346] D. A. Thorpe. The quartercentenary model of D-N explanation. Philosophy of Science, 41(2):188–195, June 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186868.

Pietarinen:1974:IIL

[347] Juhani Pietarinen. Inductive immodesty and lawlikeness. Philosophy of Science, 41(2):196–198, June 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186869.

Cartwright:1974:VFM

[348] Nancy Cartwright. Van Fraassen’s modal model of quantum mechanics. Philosophy of Science, 41(2):199–202, June 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186870.

Bradie:1974:PMP

[349] Michael Bradie. Polanyi on the Meno Paradox. Philosophy of Science, 41(2):203, June 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186871. See

comments [453].

Dale:1974:PCP

[350] A. I. Dale. On a problem in conditional probability. Philosophy of Science, 41(2):204–206, June 1974. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186872.

Ashworth:1974:BRBb

[351] E. J. Ashworth. Book review: Science and Religion in Seventeenth Century England by Richard S. Westfall. Philosophy of Science, 41(2):207, June 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/186873.

Cartwright:1974:BRB

[352] Nancy Cartwright. Book review: Paradigms and Paradoxes: The Philosophical Challenge of the Quantum Domain by Robert G. Colodny. Philosophy of Science, 41(2):207–209, June 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186874.

Anonymous:1974:BMb

[353] Anonymous. Back matter. Philosophy of Science, 41(2):210, June 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186875.

Anonymous:1974:FMb

[354] Anonymous. Front matter. Philosophy of Science, 41(2):??, June 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186863.

Spielman:1974:LTS

[355] Stephen Spielman. The logic of tests of significance. Philosophy of Science, 41(3):211–226, September 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187132.

Hellman:1974:NRR

[356] Geoffrey Hellman. The new riddle of radical translation. Philosophy of Science, 41(3):227–246, September 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187133.

Maxwell:1974:RSDb

[357] Nicholas Maxwell. The rationality of scientific discovery Part II: An aim oriented theory of scientific discovery. Philosophy of Science, 41(3):247– 295, September 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187134.

George:1974:BRB

[358] Rolf A. George. Book review: Selective History of Theories of Visual Perception: 1650–1950 by Nicholas Pastore. Philosophy of Science, 41(3): 296–297, September 1974. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187135.

Anonymous:1974:RB

[359] Anonymous. Recent books. Philosophy of Science, 41(3):298, September 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187137.

Michalos:1974:BRB

[360] Alex C. Michalos. Book review: The Coherence Theory of Truth by Nicholas Rescher. Philosophy of Science, 41(3):298, September 1974. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187136.

Anonymous:1974:BMc

[361] Anonymous. Back matter. Philosophy of Science, 41(3):??, September 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187138.

Anonymous:1974:FMc

[362] Anonymous. Front matter. Philosophy of Science, 41(3):??, September 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187131.

Clarke:1974:QTC

[363] C. J. S. Clarke. Quantum theory and cosmolog. Philosophy of Science, 41 (4):317–332, December 1974. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187002.

Grossman:1974:IID

[364] Neal Grossman. The ignorance interpretation defended. Philosophy of Science, 41(4):333–344, December 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187003.

Wright:1974:MPB

[365] Larry Wright. Mechanisms and purposive behavior III. Philosophy of Science, 41(4):345–360, December 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187004.

Holman:1974:EMO

[366] Eric W. Holman. Extensive measurement without an order relation. Philosophy of Science, 41(4):361–373, December 1974. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187005.

Narens:1974:MAA

[367] Louis Narens. Measurement without Archimedean axioms. Philosophy of Science, 41(4):374–393, December 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187006.

Maund:1974:CTR

[368] B. Maund. The conventionality of temporal relations in Relativity theory. Philosophy of Science, 41(4):394–407, December 1974. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187007.

Wilks:1974:MFD

[369] Yorick Wilks. More on Fodor’s distinction between strong and weak simulations. Philosophy of Science, 41(4):408–411, December 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187008.

Rosenberg:1974:SWF

[370] A. Rosenberg and N. J. Mackintosh. Strong, weak, and functional equivalence in machine simulation. Philosophy of Science, 41(4):412–414, December 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187009.

Trenn:1974:BRB

[371] Thaddeus J. Trenn. Book review: Thematic Origins of Scientific Thought: Kepler to Einstein by Gerald Holton. Philosophy of Science, 41(4):415– 418, December 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187010.

Berlinski:1974:BRB

[372] David Berlinski. Book review: The Philosophy of Biology by Michael Ruse. Philosophy of Science, 41(4):418–422, December 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187011.

Nolan:1974:BRB

[373] Rita Nolan. Book review: Logic Matters by P. T. Geach. Philosophy of Science, 41(4):422–424, December 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187012.

Boland:1974:BRB

[374] L. A. Boland. Book review: Epistemics & Economics: A Critique of Economic Doctrines by G. L. S. Shackle. Philosophy of Science, 41(4): 424–426, December 1974. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187013.

Krausz:1974:BRB

[375] Michael Krausz. Book review: Karl Popper by Bryan Magee. Philosophy of Science, 41(4):426–427, December 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187014.

Morgan:1974:BRBb

[376] Charles G. Morgan. Book review: Likelihood: An Account of the Statistical Concept of Likelihood and Its Application to Scientific Inference by A. W. F. Edwards. Philosophy of Science, 41(4):427–429, December 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187015.

Tuomela:1974:BRB

[377] Raimo Tuomela. Book review: Method, Model and Matter by Mario Bunge. Philosophy of Science, 41(4):429–430, December 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187016.

Rosenberg:1974:BRB

[378] Alexander Rosenberg. Book review: History as Applied Science: A Philosophical Study by William Todd. Philosophy of Science, 41(4):430–431, December 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187017.

Bub:1974:BRB

[379] Jeffrey Bub. Book review: The Conceptual Foundations of Contemporary Relativity Theory by J. C. Graves. Philosophy of Science, 41(4):431–433, December 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187018.

Anonymous:1974:AI

[380] Anonymous. Abstracts from inquiry. Philosophy of Science, 41(4):433– 434, December 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187019.

Anonymous:1974:VI

[381] Anonymous. Volume information. Philosophy of Science, 41(4):435–437, December 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187000.

Anonymous:1974:BMd

[382] Anonymous. Back matter. Philosophy of Science, 41(4):438, December 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187020.

Anonymous:1974:FMd

[383] Anonymous. Front matter. Philosophy of Science, 41(4):??, December 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187001.

Rosen:1975:ALN

[384] Deborah A. Rosen. An argument for the logical notion of a memory trace. Philosophy of Science, 42(1):1–10, March 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187291.

Hoy:1975:RGC

[385] Ronald C. Hoy. The role of genidentity in the causal theory of time. Philosophy of Science, 42(1):11–19, March 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187292.

Darmstadter:1975:BT

[386] Howard Darmstadter. Better theories. Philosophy of Science, 42(1):20– 27, March 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187293.

Hendry:1975:RSI

[387] Herbert E. Hendry. Ramsey sentences for infinite theories. Philosophy of Science, 42(1):28, March 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187294.

Yoshida:1975:FDT

[388] R. M. Yoshida. Five Duhemian theses. Philosophy of Science, 42(1):29– 45, March 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187295.

Stemmer:1975:RNN

[389] Nathan Stemmer. A relative notion of natural generalization. Philosophy of Science, 42(1):46–48, March 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187296.

McEvoy:1975:RPS

[390] John G. McEvoy. A “revolutionary” philosophy of science: Feyerabend and the degeneration of critical rationalism into sceptical fallibilism. Philosophy of Science, 42(1):49–66, March 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187297.

MacKinnon:1975:RHC

[391] Edward MacKinnon. A reinterpretation of Harre’s Copernican revolution. Philosophy of Science, 42(1):67–79, March 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187298.

Tichy:1975:WDW

[392] Pavel Tichy. What do we talk about? Philosophy of Science, 42(1):80– 93, March 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187299.

Kirschenmann:1975:BRB

[393] Peter Kirschenmann. Book review: Philosophy of Physics by Mario Bunge. Philosophy of Science, 42(1):94–96, March 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187300.

MacKinnon:1975:BRB

[394] Edward MacKinnon. Book review: The Logical Analysis of Quantum Mechanics by Erhard Scheibe; J. B. Sykes. Philosophy of Science, 42(1): 96–100, March 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187301.

Richards:1975:BRB

[395] John Richards and John H. Pfitsch. Book review: Temporal Logic by Nicholas Rescher; Alasdair Urquhart. Philosophy of Science, 42(1):100– 103, March 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187302.

Simon:1975:BRB

[396] Thomas W. Simon. Book reviews: Purposive Explanation in Psychology by Margaret A. Boden; Behavior: The Control of Perception by William T. Powers. Philosophy of Science, 42(1):103–106, March 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187303.

Hull:1975:BRB

[397] Richard T. Hull. Book review: The Conduct of Science by Michael W. Friedlander. Philosophy of Science, 42(1):106, March 1975. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187304.

Hooker:1975:BRBa

[398] C. A. Hooker. Book review: The Structure of Scientific Theories by Frederick R. Suppe. Philosophy of Science, 42(1):107, March 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187305.

Anonymous:1975:AIa

[399] Anonymous. Abstracts from inquiry. Philosophy of Science, 42(1):108– 112, March 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187306.

Anonymous:1975:RBa

[400] Anonymous. Recent books. Philosophy of Science, 42(1):112, March 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187307.

Anonymous:1975:BMa

[401] Anonymous. Back matter. Philosophy of Science, 42(1):113–136, March 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187308.

Anonymous:1975:FMa

[402] Anonymous. Front matter. Philosophy of Science, 42(1):??, March 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187290.

Falmagne:1975:SIA

[403] Jean-Claude Falmagne. A set of independent axioms for positive holder systems. Philosophy of Science, 42(2):137–151, June 1975. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187365.

Hooker:1975:GT

[404] C. A. Hooker. On global theories. Philosophy of Science, 42(2):152– 179, June 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187366.

Wendell-Waechtler:1975:MPA

[405] S. Wendell-Waechtler and E. Levy. More philosophical aspects of molecular biology. Philosophy of Science, 42(2):180–186, June 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187367.

Nowak:1975:RTC

[406] Leszek Nowak. Relative truth, the correspondence principle and absolute truth. Philosophy of Science, 42(2):187–202, June 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187368.

Ostien:1975:NDO

[407] Philip A. Ostien. A note on a definition of ’observation term’. Philosophy of Science, 42(2):203–207, June 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187369.

Angelmar:1975:MM

[408] Reinhard Angelmar and Christian Pinson. The meaning of ’marketing’. Philosophy of Science, 42(2):208–214, June 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187370.

Anonymous:1975:AIb

[409] Anonymous. Abstracts from inquiry. Philosophy of Science, 42(2):215– 221, June 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187371.

Anonymous:1975:BMb

[410] Anonymous. Back matter. Philosophy of Science, 42(2):??, June 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187372.

Anonymous:1975:FMb

[411] Anonymous. Front matter. Philosophy of Science, 42(2):??, June 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187364.

Falk:1975:LRO

[412] Arthur E. Falk. Learning to report One’s introspections. Philosophy of Science, 42(3):223–241, September 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187195.

Blair:1975:PPT

[413] David G. Blair. On purely probabilistic theories of scientific inference. Philosophy of Science, 42(3):242–249, September 1975. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187196.

Churchland:1975:TGE

[414] Paul M. Churchland. Two grades of evidential bias. Philosophy of Science, 42(3):250–259, September 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187197.

Pinkham:1975:SDA

[415] Gordon N. Pinkham. Some doubts about scientific data. Philosophy of Science, 42(3):260–269, September 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187198.

Koertge:1975:ESR

[416] Noretta Koertge. An exploration of Salmon’s S-R model of explanation. Philosophy of Science, 42(3):270–274, September 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187199.

Margolis:1975:PBM

[417] Joseph Margolis. Puccetti on brains, minds, and persons. Philosophy of Science, 42(3):275–280, September 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187200.

Puccetti:1975:RPM

[418] Roland Puccetti. A reply to Professor Margolis. Philosophy of Science, 42 (3):281–285, September 1975. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187201.

Goosens:1975:DTO

[419] William K. Goosens. Duhem’s thesis, observationality, and justification. Philosophy of Science, 42(3):286–298, September 1975. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187202.

Minton:1975:WTE

[420] Arthur J. Minton. Wright and taylor: Empiricist teleology. Philosophy of Science, 42(3):299–306, September 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187203.

Machan:1975:BBR

[421] Tibor R. Machan and M. L. Zupan. Back to being reasonable. Philosophy of Science, 42(3):307–310, September 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/ stable/187204. See response [422].

Schagrin:1975:RMZ

[422] Morton Schagrin. A response to Machan and Zupan. Philosophy of Science, 42(3):311, September 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 187205. See [421].

Goldstein:1975:NHI

[423] Leon J. Goldstein. A note on historical interpretation. Philosophy of Science, 42(3):312–319, September 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187206.

Fetzer:1975:BRA

[424] James H. Fetzer. Book review: Achinstein’s law and explanation: Law and Explanation: An Essay in the Philosophy of Science. Philosophy of Science, 42(3):320–333, September 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187207.

Hooker:1975:BRBb

[425] C. A. Hooker. Book review: Philosophical Problems of Space and Time by Adolf Grünbaum. Philosophy of Science, 42(3):334–335, September 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187208.

Brand:1975:BRB

[426] Myles Brand. Book review: The Cement of the Universe: A Study of Causation by J. L. Mackie. Philosophy of Science, 42(3):335–337, September 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187209.

Radner:1975:BRB

[427] Michael Radner. Book review: The Development of Bertrand Russell’s Philosophy by Ronald Jager. Philosophy of Science, 42(3):337, September 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187210.

Ruse:1975:BRB

[428] Michael Ruse. Book review: Darwin and His Critics: The Reception of Darwin’s Theory of Evolution by the Scientific Community by David Hull. Philosophy of Science, 42(3):338–339, September 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187211.

Wilson:1975:BRB

[429] Fred Wilson. Book review: Ontology and the Vicious-Circle Principle by Charles S. Chihara. Philosophy of Science, 42(3):339–341, September 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187212.

Creary:1975:BRB

[430] Lewis G. Creary and Christopher S. Hill. Book review: Counterfactuals by David Lewis. Philosophy of Science, 42(3):341–344, September 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187213.

Anonymous:1975:RBb

[431] Anonymous. Recent books. Philosophy of Science, 42(3):345, September 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187214.

Anonymous:1975:BMc

[432] Anonymous. Back matter. Philosophy of Science, 42(3):347–348, September 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/187215.

Anonymous:1975:FMc

[433] Anonymous. Front matter. Philosophy of Science, 42(3):??, September 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187194.

Gaa:1975:RST

[434] James Gaa. The replacement of scientific theories: Reduction and explication. Philosophy of Science, 42(4):349–372, December 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187218.

Grier:1975:PET

[435] Brown Grier. Prediction, explanation, and testability as criteria for judging statistical theories. Philosophy of Science, 42(4):373–383, December 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187219.

Stroll:1975:TT

[436] Avrum Stroll and Henry Alexander. ‘true’ and truth. Philosophy of Science, 42(4):384–410, December 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187220.

Hockney:1975:BST

[437] Donald Hockney. The bifurcation of scientific theories and indeterminacy of translation. Philosophy of Science, 42(4):411–427, December 1975. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187221.

Munson:1975:BPS

[438] Ronald Munson. Is biology a provincial science? Philosophy of Science, 42 (4):428–447, December 1975. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187222.

Marchal:1975:CS

[439] J. H. Marchal. On the concept of a system. Philosophy of Science, 42 (4):448–468, December 1975. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187223.

Baublys:1975:CSR

[440] Kenneth K. Baublys. Comments on some recent analyses of functional statements in biology. Philosophy of Science, 42(4):469–486, December 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187224.

Swoyer:1975:TCP

[441] Chris Swoyer and Thomas C. Monson. Theory confirmation in psychology. Philosophy of Science, 42(4):487–502, December 1975. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187225.

Schulman:1975:MIS

[442] L. S. Schulman, R. G. Newton, and R. Shtokhamer. Model of implication in statistical mechanics. Philosophy of Science, 42(4):503–511, December 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187226.

Anonymous:1975:VI

[443] Anonymous. Volume information. Philosophy of Science, 42(4):512–514, December 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187216.

Anonymous:1975:BMd

[444] Anonymous. Back matter. Philosophy of Science, 42(4):??, December 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187227.

Anonymous:1975:FMd

[445] Anonymous. Front matter. Philosophy of Science, 42(4):??, December 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187217.

Sklar:1976:IGM

[446] Lawrence Sklar. Inertia, gravitation and metaphysics. Philosophy of Science, 43(1):1–23, March 1976. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187333.

Nelson:1976:MR

[447] R. J. Nelson. On mechanical recognition. Philosophy of Science, 43(1):24– 52, March 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187334.

Allen:1976:NPU

[448] Edward H. Allen. Negative probabilities and the uses of signed probability theory. Philosophy of Science, 43(1):53–70, March 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187335.

Rescher:1976:PER

[449] Nicholas Rescher. Peirce and the economy of research. Philosophy of Science, 43(1):71–98, March 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187336.

Kitcher:1976:HE

[450] Philip Kitcher. Hilbert’s epistemology. Philosophy of Science, 43(1):99– 115, March 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187337.

Carlson:1976:LTS

[451] Roger Carlson. The logic of tests of significance. Philosophy of Science, 43(1):116–128, March 1976. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187338.

Kruger:1976:SEP

[452] Lorenz Kruger. Are statistical explanations possible? Philosophy of Science, 43(1):129–146, March 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187339.

Simon:1976:BPM

[453] Herbert A. Simon. Bradie on Polanyi on the Meno Paradox. Philosophy of Science, 43(1):147–150, March 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 187340. See [349].

Friedman:1976:PVE

[454] K. S. Friedman. A partial vindication of ergodic theory. Philosophy of Science, 43(1):151–162, March 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187341.

Anonymous:1976:BMa

[455] Anonymous. Back matter. Philosophy of Science, 43(1):163–164, March 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187342.

Anonymous:1976:FMa

[456] Anonymous. Front matter. Philosophy of Science, 43(1):??, March 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187332.

Grunbaum:1976:PPD

[457] Adolf Grünbaum. Is preacceleration of particles in Dirac’s electrodynamics a case of backward causation? The myth of retrocausation in classical electrodynamics. Philosophy of Science, 43(2):165–201, June 1976. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187262. See corrections [522].

Cocchiarella:1976:LNK

[458] Nino Cocchiarella. On the logic of natural kinds. Philosophy of Science, 43(2):202–222, June 1976. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187263.

Ringen:1976:ETO

[459] Jon D. Ringen. Explanation, teleology, and operant behaviorism: A study of the experimental analysis of purposive behavior. Philosophy of Science, 43(2):223–253, June 1976. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187264.

Dale:1976:PL

[460] A. I. Dale. Probability logic and F. Philosophy of Science, 43(2):254– 265, June 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187265.

Levin:1976:ECC

[461] Michael E. Levin. The extensionality of causation and causal-explanatory contexts. Philosophy of Science, 43(2):266–277, June 1976. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187266.

Vetterling:1976:MRP

[462] Mary K. Vetterling. More on reflexive predictions. Philosophy of Science, 43(2):278–282, June 1976. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187267.

Faber:1976:REC

[463] Roger J. Faber. Re-encountering a counter-intuitive probability. Philosophy of Science, 43(2):283–285, June 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187268.

Goldberg:1976:CCP

[464] Samuel Goldberg. Copi’s conditional probability problem. Philosophy of Science, 43(2):286–289, June 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187269.

Hesslow:1976:TNP

[465] Germund Hesslow. Two notes on the probabilistic approach to causality. Philosophy of Science, 43(2):290–292, June 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187270.

Kockelmans:1976:BRB

[466] Joseph J. Kockelmans. Book review: Theorie und Erfahrung. Zweiter Halbband: Theorienstrukturen und Theoriendynamik by Wolfgang Stegmuller. Philosophy of Science, 43(2):293–297, June 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http:/

/www.jstor.org/stable/187271.

Margolis:1976:BRB

[467] Joseph Margolis. Book review: Logic and Art. Essays in Honor of Nelson Goodman by Richard Rudner; Israel Scheffler. Philosophy of Science, 43 (2):297–301, June 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187272.

Corcoran:1976:BRB

[468] John Corcoran and Stewart Shapiro. Book review: What Is Mathematical Logic? by John N. Crossley. Philosophy of Science, 43(2):301–302, June 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187273.

Ruse:1976:BRB

[469] Michael Ruse. Book review: The Genetic Basis of Evolutionary Change by R. C. Lewontin. Philosophy of Science, 43(2):302–304, June 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187274.

Caplan:1976:BRB

[470] Arthur Caplan. Book review: Sociobiology by Edward O. Wilson. Philosophy of Science, 43(2):305–306, June 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187275.

Anonymous:1976:A

[471] Anonymous. Announcement. Philosophy of Science, 43(2):307, June 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187276.

Anonymous:1976:BMb

[472] Anonymous. Back matter. Philosophy of Science, 43(2):308–309, June 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187277.

Anonymous:1976:FMb

[473] Anonymous. Front matter. Philosophy of Science, 43(2):??, June 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187261.

Keuth:1976:VAW

[474] Herbert Keuth. Verisimilitude or the approach to the whole truth. Philosophy of Science, 43(3):311–336, September 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187229.

McCall:1976:OTF

[475] Storrs McCall. Objective time flow. Philosophy of Science, 43(3):337–362, September 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187230.

Moreland:1976:PG

[476] John Moreland. On projecting grue. Philosophy of Science, 43(3):363–377, September 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187231.

Brown:1976:SCM

[477] Cecil H. Brown. Semantic components, meaning, and use in ethnosemantics. Philosophy of Science, 43(3):378–395, September 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187232.

Quinn:1976:IMC

[478] Philip L. Quinn. Intrinsic metrics on continuous spatial manifolds. Philosophy of Science, 43(3):396–414, September 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187233.

Brown:1976:CEU

[479] Peter M. Brown. Conditionalization and expected utility. Philosophy of Science, 43(3):415–419, September 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187234.

Gardenfors:1976:RRD

[480] Peter Gardenfors. Relevance and redundancy in deductive explanations. Philosophy of Science, 43(3):420–431, September 1976. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187235.

Priest:1976:GS

[481] Graham Priest. Gruesome simplicity. Philosophy of Science, 43(3):432– 437, September 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187236.

Rolston:1976:NSP

[482] Howard L. Rolston. A note of simplicity as a principle for evaluating rival scientific theories. Philosophy of Science, 43(3):438–440, September 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187237.

Siegel:1976:MSK

[483] Harvey Siegel. Meiland on Scheffler, Kuhn, and objectivity in science. Philosophy of Science, 43(3):441–448, September 1976. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187238.

Lugg:1976:BRB

[484] Andrew Lugg. Book review: The Interpretation of Quantum Mechanics by Michael Audi. Philosophy of Science, 43(3):449–452, September 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187239.

Burian:1976:BRB

[485] Richard M. Burian. Book review: Theoretical Concepts by Raimo Tuomela. Philosophy of Science, 43(3):452–456, September 1976. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187240.

Simon:1976:BRB

[486] Thomas W. Simon. Book review: On Purposeful Systems by Russell L. Ackoff; Fred E. Emery. Philosophy of Science, 43(3):456–458, September 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187241.

Watson:1976:BRB

[487] Richard A. Watson. Book reviews: Recent Earth History by Claudio Vita-Finzi; The Nature of the Stratigraphical Record by Derek V. Ager. Philosophy of Science, 43(3):458–459, September 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187242.

Anonymous:1976:AI

[488] Anonymous. Abstracts from inquiry. Philosophy of Science, 43(3):459– 467, September 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187243.

Anonymous:1976:BMc

[489] Anonymous. Back matter. Philosophy of Science, 43(3):??, September 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187244.

Anonymous:1976:FMc

[490] Anonymous. Front matter. Philosophy of Science, 43(3):??, September 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187228.

Beauregard:1976:SGC

[491] Laurent A. Beauregard. The sui generis conventionality of simultaneity. Philosophy of Science, 43(4):469–490, December 1976. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187114.

Cortes:1976:LPI

[492] Alberto Cortes. Leibniz’s Principle of the Identity of Indiscernibles: A false principle. Philosophy of Science, 43(4):491–505, December 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187115.

Derden:1976:CDA

[493] J. K. Derden, Jr. Carnap’s definition of ‘analytic truth’ for scientific theories. Philosophy of Science, 43(4):506–522, December 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187116.

Fumerton:1976:SC

[494] R. A. Fumerton. Subjunctive conditionals. Philosophy of Science, 43(4): 523–538, December 1976. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187117.

VanValen:1976:IC

[495] Leigh Van Valen. Individualistic classes. Philosophy of Science, 43(4): 539–541, December 1976. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187118.

Jobe:1976:PCD

[496] Evan K. Jobe. A puzzle concerning D-N explanation. Philosophy of Science, 43(4):542–549, December 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187119.

Moor:1976:CSC

[497] James H. Moor. The cancellation of symmetrical contraries and the principle of significant contradictories. Philosophy of Science, 43(4):550–559, December 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187120.

Frankel:1976:HC

[498] Henry Frankel. Harre on causation. Philosophy of Science, 43(4):560–569, December 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187121.

Girill:1976:MEP

[499] T. R. Girill. Are micro-entities picturable? Philosophy of Science, 43 (4):570–574, December 1976. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187122.

Swijtink:1976:TSR

[500] Zeno G. Swijtink. Two suggestions for Ramsey-reducts of infinite theories. Philosophy of Science, 43(4):575–577, December 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187123.

Kerr:1976:MWB

[501] Stanley Kerr. Many worlds are better than none. Philosophy of Science, 43 (4):578–582, December 1976. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187124.

Clarke:1976:RSK

[502] C. J. S. Clarke. Reply to Stanley Kerr. Philosophy of Science, 43(4): 583–584, December 1976. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187125.

Laudan:1976:TDM

[503] Larry Laudan. Two dogmas of methodology. Philosophy of Science, 43 (4):585–597, December 1976. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187126.

Rietdijk:1976:SRD

[504] C. W. Rietdijk. Special Relativity and determinism. Philosophy of Science, 43(4):598–609, December 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187127.

Hoy:1976:NGB

[505] Ronald C. Hoy. A note on Gustav Bergmann’s treatment of temporal consciousness. Philosophy of Science, 43(4):610–617, December 1976. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187128.

Fitzgerald:1976:BRS

[506] Paul Fitzgerald. Book review: Swinburne’s Space and Time. Philosophy of Science, 43(4):618–637, December 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187129.

Anonymous:1976:VI

[507] Anonymous. Volume information. Philosophy of Science, 43(4):639–644, December 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187112.

Anonymous:1976:BMd

[508] Anonymous. Back matter. Philosophy of Science, 43(4):645–646, December 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187130.

Anonymous:1976:FMd

[509] Anonymous. Front matter. Philosophy of Science, 43(4):??, December 1976. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187113.

Burian:1977:MTM

[510] Richard M. Burian. More than a marriage of convenience: On the inextricability of history and philosophy of science. Philosophy of Science, 44(1): 1–42, March 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187098.

Darden:1977:IT

[511] Lindley Darden and Nancy Maull. Interfield theories. Philosophy of Science, 44(1):43–64, March 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187099.

Norton:1977:MNC

[512] Bryan G. Norton. On the metatheoretical nature of Carnap’s philosophy. Philosophy of Science, 44(1):65–85, March 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187100.

Glymour:1977:SRW

[513] Clark Glymour. The sum rule is well-confirmed. Philosophy of Science, 44(1):86–94, March 1977. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187101.

Fine:1977:CSR

[514] Arthur Fine. Conservation, the sum rule and confirmation. Philosophy of Science, 44(1):95–106, March 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187102.

Cartwright:1977:SRB

[515] Nancy Cartwright. The sum rule has not been tested. Philosophy of Science, 44(1):107–112, March 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187103.

Batchelder:1977:CEA

[516] William H. Batchelder and Louis Narens. A critical examination of the analysis of dichotomous data. Philosophy of Science, 44(1):113–135, March 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/187104.

Shrader:1977:CES

[517] Douglas W. Shrader, Jr. Causation, explanation, and statistical relevance. Philosophy of Science, 44(1):136–145, March 1977. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187105.

Harre:1977:SRR

[518] Rom Harre. Science as representation: A reply to Mr. MacKinnon. Philosophy of Science, 44(1):146–158, March 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187106.

McCarthy:1977:AMS

[519] Timothy McCarthy. On an Aristotelian model of scientific explanation. Philosophy of Science, 44(1):159–166, March 1977. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187107.

Weingard:1977:BRB

[520] Robert Weingard. Book review: Space, Time and Spacetime by Lawrence Sklar. Philosophy of Science, 44(1):167–173, March 1977. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187108.

Adler:1977:BRB

[521] Jonathan E. Adler. Book review: Local Induction by Radu J. Bogdan. Philosophy of Science, 44(1):173–177, March 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187109.

Anonymous:1977:CPP

[522] Anonymous. Corrections: Is Preacceleration of Particles in Dirac’s Electrodynamics a Case of Backward Causation? The Myth of Retrocausation in Classical Electrodynamics. Philosophy of Science, 44(1):177, March 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187110. See [457].

Anonymous:1977:BMa

[523] Anonymous. Back matter. Philosophy of Science, 44(1):??, March 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187111.

Anonymous:1977:FMa

[524] Anonymous. Front matter. Philosophy of Science, 44(1):??, March 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187097.

Salmon:1977:HCI

[525] Wesley C. Salmon. Hempel’s conception of inductive inference in inductive-statistical explanation. Philosophy of Science, 44(2):179–185, June 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/187344.

Coffa:1977:PRT

[526] J. Alberto Coffa. Probabilities: Reasonable or true? Philosophy of Science, 44(2):186–198, June 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187345.

Salmon:1977:IER

[527] Wesley C. Salmon. Indeterminism and epistemic relativization. Philosophy of Science, 44(2):199–202, June 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187346.

Sayre:1977:SMC

[528] Kenneth M. Sayre. Statistical models of causal relations. Philosophy of Science, 44(2):203–214, June 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187347.

Salmon:1977:TCI

[529] Wesley C. Salmon. An “at-at” theory of causal influence. Philosophy of Science, 44(2):215–224, June 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187348.

Laymon:1977:FBM

[530] Ronald Laymon. Feyerabend, Brownian motion, and the hiddenness of refuting facts. Philosophy of Science, 44(2):225–247, June 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187349.

Dretske:1977:LN

[531] Fred I. Dretske. Laws of nature. Philosophy of Science, 44(2):248–268, June 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/187350.

Cummins:1977:PEB

[532] Robert Cummins. Programs in the explanation of behavior. Philosophy of Science, 44(2):269–287, June 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187351.

Weingard:1977:CNA

[533] Robert Weingard. On cracking that nut, absolute space. Philosophy of Science, 44(2):288–291, June 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187352.

Meyer:1977:URH

[534] Stuart L. Meyer. Urning a resolution of Hempel’s paradox. Philosophy of Science, 44(2):292–296, June 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187353.

Pastore:1977:RGT

[535] Nicholas Pastore. Reply to George: Thomas Reid and the constancy hypothesis. Philosophy of Science, 44(2):297–302, June 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187354.

Puccetti:1977:GCF

[536] Roland Puccetti. The great C-fiber myth: A critical note. Philosophy of Science, 44(2):303–305, June 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187355.

Schlegel:1977:CPS

[537] Richard Schlegel. The clock paradox: Some new thoughts. Philosophy of Science, 44(2):306–312, June 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187356.

Utz:1977:TO

[538] Stephen Utz. On teleology and organisms. Philosophy of Science, 44(2): 313–320, June 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187357.

Wright:1977:RU

[539] Larry Wright. Rejoinder to utz. Philosophy of Science, 44(2):321–325, June 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/187358.

Langtry:1977:PII

[540] Bruce Langtry. Popper on induction and Independence. Philosophy of Science, 44(2):326–331, June 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187359.

Sklar:1977:BRB

[541] Lawrence Sklar. Book review: The Philosophy of Quantum Mechanics: The Interpretations of Quantum Mechanics in Historical Perspective by Max Jammer. Philosophy of Science, 44(2):332, June 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187360.

Levy:1977:BRB

[542] E. Levy, A. Chernavska, and R. I. G. Hughes. Book review: The Interpretation of Quantum Mechanics by Jeffrey Bub. Philosophy of Science, 44(2):332–336, June 1977. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187361.

Rosenthal:1977:BRB

[543] Sandra B. Rosenthal. Book review: Four Pragmatists: A Critical Introduction to Peirce, James, Mead, and Dewey by Israel Scheffler. Philosophy of Science, 44(2):336–339, June 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187362.

Anonymous:1977:BMb

[544] Anonymous. Back matter. Philosophy of Science, 44(2):340, June 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187363.

Anonymous:1977:FMb

[545] Anonymous. Front matter. Philosophy of Science, 44(2):??, June 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187343.

Achinstein:1977:FS

[546] Peter Achinstein. Function statements. Philosophy of Science, 44(3): 341–367, September 1977. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187388.

Levi:1977:EUE

[547] Isaac Levi. Epistemic utility and the evaluation of experiments. Philosophy of Science, 44(3):368–386, September 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187389.

Cupples:1977:TTE

[548] Brian Cupples. Three types of explanation. Philosophy of Science, 44(3): 387–408, September 1977. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187390.

Shrader-Frechette:1977:ACA

[549] K. Shrader-Frechette. Atomism in crisis: An analysis of the current high energy paradigm. Philosophy of Science, 44(3):409–440, September 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187391.

Bradie:1977:DRS

[550] Michael P. Bradie. The development of Russell’s structural postulates. Philosophy of Science, 44(3):441–463, September 1977. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187392.

Jackson:1977:RSS

[551] Frank Jackson and Robert Pargetter. Relative simultaneity in the Special Relativity. Philosophy of Science, 44(3):464–474, September 1977. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187393. See comment [677].

Teller:1977:PHV

[552] Paul Teller. On the problem of hidden variables for quantum mechanical observables with continuous spectra. Philosophy of Science, 44(3):475– 477, September 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187394.

Oaklander:1977:PPQ

[553] L. Nathan Oaklander. Particulars, positional qualities, and individuation. Philosophy of Science, 44(3):478–490, September 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187395. See comments [554].

Bergmann:1977:SCP

[554] Gustav Bergmann. Some comments on Professor Oaklander’s “Particulars, Positional Qualities, and Individuation”. Philosophy of Science, 44 (3):491–493, September 1977. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187396.

See [553].

Richards:1977:NSM

[555] Robert J. Richards. The natural selection model of conceptual evolution. Philosophy of Science, 44(3):494–501, September 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187397. See comment [556].

Campbell:1977:CBN

[556] Donald T. Campbell. Comment on “The Natural Selection Model of Conceptual Evolution”. Philosophy of Science, 44(3):502–507, September 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187398.

Anonymous:1977:A

[557] Anonymous. Announcement. Philosophy of Science, 44(3):508, September 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187399.

Anonymous:1977:BMc

[558] Anonymous. Back matter. Philosophy of Science, 44(3):??, September 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187400.

Anonymous:1977:FMc

[559] Anonymous. Front matter. Philosophy of Science, 44(3):??, September 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187387.

Anonymous:1977:EIS

[560] Anonymous. Editorial introduction to the special section on value issues in science, technology, and medicine. Philosophy of Science, 44(4):511–512, December 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186937.

Gaa:1977:MAR

[561] James C. Gaa. Moral autonomy and the rationality of science. Philosophy of Science, 44(4):513–541, December 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186938.

Boorse:1977:HTC

[562] Christopher Boorse. Health as a theoretical concept. Philosophy of Science, 44(4):542–573, December 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186939.

Carpenter:1977:PIT

[563] Stanley R. Carpenter. Philosophical issues in technology assessment. Philosophy of Science, 44(4):574–593, December 1977. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186940.

Singer:1977:JTT

[564] Marcus G. Singer. Justice, theory, and a theory of justice. Philosophy of Science, 44(4):594–618, December 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186941.

Whitbeck:1977:CMD

[565] Caroline Whitbeck. Causation in medicine: The disease entity model. Philosophy of Science, 44(4):619–637, December 1977. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186942.

Ruse:1977:KPP

[566] Michael Ruse. Karl Popper’s philosophy of biology. Philosophy of Science, 44(4):638–661, December 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186943.

Hull:1977:BRB

[567] David L. Hull. Book review: The Collected Papers of Charles Darwin by Paul H. Barrett. Philosophy of Science, 44(4):662–663, December 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186944.

Green:1977:BRB

[568] Edward J. Green. Book review: Method and Appraisal in Economics by Sprio Latsis. Philosophy of Science, 44(4):663–666, December 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186945.

MacKay:1977:BRB

[569] Alfred F. MacKay. Book review: Microeconomic Laws: A Philosophical Analysis by Alexander Rosenberg. Philosophy of Science, 44(4):666–668, December 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186946.

Anonymous:1977:VI

[570] Anonymous. Volume information. Philosophy of Science, 44(4):669–673, December 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186935.

Anonymous:1977:BMd

[571] Anonymous. Back matter. Philosophy of Science, 44(4):??, December 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186947.

Anonymous:1977:FMd

[572] Anonymous. Front matter. Philosophy of Science, 44(4):??, December 1977. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186936.

Luce:1978:DIN

[573] R. Duncan Luce. Dimensionally invariant numerical laws correspond to meaningful qualitative relations. Philosophy of Science, 45(1):1–16, March 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186895.

Giannoni:1978:RME

[574] Carlo Giannoni. Relativistic mechanics and electrodynamics without oneway velocity assumptions. Philosophy of Science, 45(1):17–46, March 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186896.

Quay:1978:PEE

[575] S. J. Paul M. Quay. A philosophical explanation of the explanatory functions of ergodic theory. Philosophy of Science, 45(1):47–59, March 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186897.

Heil:1978:TTP

[576] John Heil. Traces of things past. Philosophy of Science, 45(1):60–72, March 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/186898.

Goosens:1978:RMG

[577] William K. Goosens. Reduction by molecular genetics. Philosophy of Science, 45(1):73–95, March 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186899.

Brandon:1978:E

[578] Robert N. Brandon. ’evolution’. Philosophy of Science, 45(1):96–109, March 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/186900.

Kordig:1978:DJ

[579] Carl R. Kordig. Discovery and justification. Philosophy of Science, 45 (1):110–117, March 1978. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186901.

Glymour:1978:TFM

[580] Clark Glymour. Two flagpoles are more paradoxical than one. Philosophy of Science, 45(1):118–119, March 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186902.

Spielman:1978:SDL

[581] Stephen Spielman. Statistical dogma and the logic of significance testing. Philosophy of Science, 45(1):120–135, March 1978. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186903.

Dorling:1978:EPS

[582] Jon Dorling. On explanations in physics: Sketch of an alternative to Hempel’s account of the explanation of laws. Philosophy of Science, 45 (1):136–140, March 1978. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186904.

Kukla:1978:ESP

[583] Andy Kukla. On the empirical significance of pure determinism. Philosophy of Science, 45(1):141–144, March 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186905.

McGrath:1978:BRB

[584] James H. McGrath. Book review: The Logico-Algebraic Approach to Quantum Mechanics. Volume I: Historical Evolution by C. A. Hooker. Philosophy of Science, 45(1):145–148, March 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/186906.

Curd:1978:BRB

[585] Martin V. Curd. Book review: Ludwig Boltzmann, Theoretical Physics and Philosophical Problems, Selected Writings by B. McGuinness; P. Foulkes. Philosophy of Science, 45(1):148–149, March 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186907.

Gilbert:1978:BRB

[586] Michael A. Gilbert. Book review: Deviant Logic: Some Philosophical Issues by Susan Haack. Philosophy of Science, 45(1):149–151, March 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186908.

Lehman:1978:BRB

[587] Hugh Lehman. Book reviews: Mathematics, Matter and Method. Philosophical Papers by Hilary Putnam; Mind, Language and Reality. Philosophical Papers by Hilary Putnam. Philosophy of Science, 45(1):151–155, March 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/186909.

Carlstrom:1978:BRB

[588] Ian F. Carlstrom and Christopher S. Hill. Book review: The Logic of Conditionals by Ernest W. Adams. Philosophy of Science, 45(1):155– 158, March 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186910.

vanFraassen:1978:BRB

[589] Bas C. van Fraassen. Book review: Personnelle and Statistische Wahrscheinlichkeit by Wolfgang Stegmuller. Philosophy of Science, 45 (1):158–163, March 1978. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186911.

Sanford:1978:BRB

[590] David H. Sanford. Book review: Perception, Common Sense, and Science by James W. Cornman. Philosophy of Science, 45(1):163–165, March 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186912.

Fitzgerald:1978:BRB

[591] Paul Fitzgerald. Book review: Perception: Facts and Theories by C. W. K. Mundle. Philosophy of Science, 45(1):165–169, March 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186913.

Anonymous:1978:BMa

[592] Anonymous. Back matter. Philosophy of Science, 45(1):170–172, March 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186914.

Anonymous:1978:FMa

[593] Anonymous. Front matter. Philosophy of Science, 45(1):??, March 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186894.

Earman:1978:UL

[594] John Earman. The universality of laws. Philosophy of Science, 45(2):173– 181, June 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186815.

Harris:1978:SST

[595] John Henry Harris. Strong scientific theories. Philosophy of Science, 45 (2):182–205, June 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186816.

Railton:1978:DNM

[596] Peter Railton. A deductive-nomological model of probabilistic explanation. Philosophy of Science, 45(2):206–226, June 1978. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186817.

Horgan:1978:SBL

[597] Terence E. Horgan. Supervenient bridge laws. Philosophy of Science, 45(2):227–249, June 1978. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186818.

Kantorovich:1978:IMG

[598] Aharon Kantorovich. An ideal model for the growth of knowledge in research programs. Philosophy of Science, 45(2):250–272, June 1978. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186819.

Fales:1978:TSD

[599] Evan Fales. Theoretical simplicity and defeasibility. Philosophy of Science, 45(2):273–288, June 1978. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186820.

Rody:1978:CIR

[600] Phillip J. Rody. (c) instances, the relevance criterion, and the paradoxes of confirmation. Philosophy of Science, 45(2):289–302, June 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186821.

Colonius:1978:WEM

[601] Hans Colonius. On weak extensive measurement. Philosophy of Science, 45(2):303–308, June 1978. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186822.

Ellis:1978:SSI

[602] Brian Ellis. Is signal synchrony independent of transport synchrony? Philosophy of Science, 45(2):309–311, June 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186823.

Schlesinger:1978:MP

[603] G. Schlesinger. On a matter of principle. Philosophy of Science, 45(2):312– 317, June 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186824.

Laymon:1978:BRB

[604] Ronald Laymon. Book review: Method and Appraisal in the Physical Sciences by Colin Howson. Philosophy of Science, 45(2):318–322, June 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186825.

Leslie:1978:BRB

[605] John Leslie. Book review: The Study of Time II by J. T. Fraser; N. Lawrence. Philosophy of Science, 45(2):322–325, June 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186826.

Burian:1978:BRB

[606] Richard M. Burian. Book review: Linguistic Representation by Jay F. Rosenberg. Philosophy of Science, 45(2):325–328, June 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186827.

McMullin:1978:BRB

[607] Ernan McMullin. Book review: Reason, Experiment and Mysticism in the Scientific Revolution by M. L. Righini Bonelli; William R. Shea. Philosophy of Science, 45(2):329–330, June 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186828.

Shea:1978:BRB

[608] William R. Shea. Book review: L’Indagine del Mondo Fisico by Giuliano Toraldo di Francia. Philosophy of Science, 45(2):330–331, June 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186829.

Dray:1978:BRB

[609] William H. Dray. Book review: History of Science as Explanation by M. A. Finocchiaro. Philosophy of Science, 45(2):331–333, June 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186830.

Anonymous:1978:BMb

[610] Anonymous. Back matter. Philosophy of Science, 45(2):334, June 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186831.

Anonymous:1978:FMb

[611] Anonymous. Front matter. Philosophy of Science, 45(2):??, June 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186814.

Hull:1978:MI

[612] David L. Hull. A matter of individuality. Philosophy of Science, 45(3): 335–360, September 1978. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187022.

Field:1978:NJC

[613] Hartry Field. A note on Jeffrey conditionalization. Philosophy of Science, 45(3):361–367, September 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187023.

Rosenberg:1978:SBC

[614] Alexander Rosenberg. The supervenience of biological concepts. Philosophy of Science, 45(3):368–386, September 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187024.

Miller:1978:MIS

[615] Richard W. Miller. Methodological individualism and social explanation. Philosophy of Science, 45(3):387–414, September 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187025.

Chao-Tien:1978:SPC

[616] Lin Chao-Tien. Solutions to the paradoxes of confirmation, Goodman’s paradox, and two new theories of confirmation. Philosophy of Science, 45 (3):415–419, September 1978. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187026.

Keuth:1978:TDT

[617] Herbert Keuth. Tarski’s definition of truth and the correspondence theory. Philosophy of Science, 45(3):420–430, September 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187027.

Niiniluoto:1978:DLN

[618] Ilkka Niiniluoto. Dretske on laws of nature. Philosophy of Science, 45(3): 431–439, September 1978. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187028.

Dretske:1978:RN

[619] Fred I. Dretske. Reply to Niiniluoto. Philosophy of Science, 45(3):440–444, September 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187029.

Minogue:1978:RIR

[620] Brendan P. Minogue. Realism and intensional reference. Philosophy of Science, 45(3):445–455, September 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187030.

Tuana:1978:QDE

[621] Nancy Tuana. Quinn on Duhem: An emendation. Philosophy of Science, 45(3):456–462, September 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187031.

Quinn:1978:RT

[622] Philip L. Quinn. Rejoinder to Tuana. Philosophy of Science, 45(3):463– 465, September 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187032.

Barnette:1978:DQM

[623] R. L. Barnette. Does quantum mechanics disprove the principle of the identity of indiscernibles? Philosophy of Science, 45(3):466–470, September 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/187033.

Lauwers:1978:RCI

[624] Steven Lauwers. A reexamination of causal irregularity. Philosophy of Science, 45(3):471–473, September 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187034.

Creath:1978:QE

[625] Richard Creath. A query on entrenchment. Philosophy of Science, 45(3): 474–477, September 1978. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187035.

Sarkar:1978:MBA

[626] Husain Sarkar. Musgrave’s “Appraisals and Advice”. Philosophy of Science, 45(3):478–483, September 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187036.

Noren:1978:MPP

[627] Stephen J. Noren. Micro-particles and picturability: A reply. Philosophy of Science, 45(3):484–487, September 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187037.

Albert:1978:BRB

[628] Daniel A. Albert and Michael D. Resnik. Book review: The Logic of Medicine by Edmond A. Murphy. Philosophy of Science, 45(3):488–490, September 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187038.

Kielkopf:1978:BRB

[629] Charles F. Kielkopf. Book review: The Logic of Questions and Answers by Nuel D. Belnap,; Thomas B. Steel. Philosophy of Science, 45(3):490–491, September 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187039.

Fetzer:1978:BRB

[630] James H. Fetzer. Book review: The Philosophy of Karl Popper by Robert John Ackermann. Philosophy of Science, 45(3):491–493, September 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187040.

Thompson:1978:BRB

[631] Manley Thompson. Book review: Methodological Pragmatism by Nicholas Rescher. Philosophy of Science, 45(3):493–495, September 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187041.

Anonymous:1978:BMc

[632] Anonymous. Back matter. Philosophy of Science, 45(3):??, September 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187042.

Anonymous:1978:FMc

[633] Anonymous. Front matter. Philosophy of Science, 45(3):??, September 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187021.

Anonymous:1978:E

[634] Anonymous. Editorial. Philosophy of Science, 45(4):497, December 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186965.

Stich:1978:BSS

[635] Stephen P. Stich. Beliefs and subdoxastic states. Philosophy of Science, 45 (4):499–518, December 1978. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186966.

Reed:1978:GTP

[636] Edward S. Reed and Rebecca K. Jones. Gibson’s theory of perception: A case of hasty epistemologizing? Philosophy of Science, 45(4):519–530, December 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186967.

Hanna:1978:TIM

[637] Joseph F. Hanna. On transmitted information as a measure of explanatory power. Philosophy of Science, 45(4):531–562, December 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186968.

Dacey:1978:TC

[638] Raymond Dacey. A theory of conclusions. Philosophy of Science, 45(4): 563–574, December 1978. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186969.

Cassidy:1978:PAG

[639] John Cassidy. Philosophical aspects of the group selection controversy. Philosophy of Science, 45(4):575–594, December 1978. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186970.

Schwartz:1978:IAP

[640] Robert John Schwartz. Idealizations and approximations in physics. Philosophy of Science, 45(4):595–603, December 1978. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186971.

Rosen:1978:DPT

[641] Deborah A. Rosen. In defense of a probabilistic theory of causality. Philosophy of Science, 45(4):604–613, December 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186972.

Stern:1978:AEC

[642] Cindy Stern. On the alleged extensionality of “causal explanatory contexts”. Philosophy of Science, 45(4):614–625, December 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186973.

Levin:1978:LSB

[643] Michael E. Levin and Margarita R. Levin. Lavoisier’s slow burn. Philosophy of Science, 45(4):626–629, December 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186974.

Shrader:1978:SSM

[644] Douglas Shrader, Jr. Sayre’s statistical model of causal relations. Philosophy of Science, 45(4):630–632, December 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186975.

Sayre:1978:MCR

[645] Kenneth Sayre. Masking and causal relatedness: An elucidation. Philosophy of Science, 45(4):633–637, December 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186976.

Heffernan:1978:SDA

[646] James D. Heffernan. Some doubts about “Turing Machine Arguments”. Philosophy of Science, 45(4):638–647, December 1978. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186977.

Dennett:1978:BRB

[647] Daniel C. Dennett. Book review: Artificial Intelligence and Natural Man by Margaret A. Boden. Philosophy of Science, 45(4):648–649, December 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186978.

Frankel:1978:BRB

[648] Henry Frankel. Book review: The Essential Tension by Thomas S. Kuhn. Philosophy of Science, 45(4):649–652, December 1978. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186979.

Hodge:1978:BRB

[649] M. J. S. Hodge. Book review: Ontogeny and Phylogeny by Stephen Jay Gould. Philosophy of Science, 45(4):652–653, December 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186980.

Levin:1978:BRB

[650] Michael E. Levin. Book review: Cybernetics and the Philosophy of Mind by Kenneth Sayre. Philosophy of Science, 45(4):653–654, December 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186981.

Anonymous:1978:VI

[651] Anonymous. Volume information. Philosophy of Science, 45(4):655–660, December 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186963.

Anonymous:1978:BMd

[652] Anonymous. Back matter. Philosophy of Science, 45(4):??, December 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186982.

Anonymous:1978:FMd

[653] Anonymous. Front matter. Philosophy of Science, 45(4):??, December 1978. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186964.

Gardner:1979:RIC

[654] Michael R. Gardner. Realism and instrumentalism in 19th-Century atomism. Philosophy of Science, 46(1):1–34, March 1979. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186800.

Longino:1979:EHA

[655] Helen E. Longino. Evidence and hypothesis: An analysis of evidential relations. Philosophy of Science, 46(1):35–56, March 1979. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186801.

Byerly:1979:SCN

[656] Henry Byerly. Substantial causes and nomic determination. Philosophy of Science, 46(1):57–81, March 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186802.

Matthews:1979:TAS

[657] Geoffrey Matthews. Time’s Arrow and the structure of spacetime. Philosophy of Science, 46(1):82–97, March 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186803.

Merrill:1979:CP

[658] G. H. Merrill. Confirmation and prediction. Philosophy of Science, 46(1): 98–117, March 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186804.

Nissim-Sabat:1979:GRC

[659] Charles Nissim-Sabat. On Grunbaum and retrocausation in classical electrodynamics. Philosophy of Science, 46(1):118–135, March 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186805. See reply [660].

Grunbaum:1979:RFA

[660] Adolf Grünbaum and Allen I. Janis. Retrocausation and the formal assimilation of classical electrodynamics to Newtonian mechanics: A reply to Nissim-Sabat’s “On Grünbaum and Retrocausation”. Philosophy of Science, 46(1):136–160, March 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 186806. See [659].

Wilson:1979:GNF

[661] Mark Wilson. Generality and nomological form. Philosophy of Science, 46(1):161–164, March 1979. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186807.

Ruse:1979:BRB

[662] Michael Ruse. Book review: The Young Darwin and His Cultural Circle by Edward Manier. Philosophy of Science, 46(1):165–166, March 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186808.

Laudan:1979:BRB

[663] Rachel Laudan. Book review: The Structure of Geology by David B. Kitts. Philosophy of Science, 46(1):166–167, March 1979. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186809.

Beaver:1979:BRB

[664] Donald de B. Beaver. Book review: Creation by Natural Law: Laplace’s Nebular Hypothesis in American Thought by Ronald L. Numbers. Philosophy of Science, 46(1):167–169, March 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186810.

Anonymous:1979:BRB

[665] Anonymous. Book review: Acts and Other Events by Judith Jarvis Thomson. Philosophy of Science, 46(1):169–170, March 1979. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186811.

Quinn:1979:BRB

[666] Philip L. Quinn. Book review: Religion and Scientific Method by George Schlesinger. Philosophy of Science, 46(1):170–171, March 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186812.

Anonymous:1979:BMa

[667] Anonymous. Back matter. Philosophy of Science, 46(1):172–174, March 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186813.

Anonymous:1979:FMa

[668] Anonymous. Front matter. Philosophy of Science, 46(1):??, March 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186799.

Rozeboom:1979:BTR

[669] William W. Rozeboom. On behavioral theories of reference. Philosophy of Science, 46(2):175–203, June 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187044.

Cyert:1979:TBM

[670] Richard M. Cyert and Garrel Pottinger. Towards a better microeconomic theory. Philosophy of Science, 46(2):204–222, June 1979. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187045.

Bigelow:1979:QPL

[671] John C. Bigelow. Quantum probability in logical space. Philosophy of Science, 46(2):223–243, June 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187046.

Moberg:1979:TRI

[672] Dale W. Moberg. Are there rival, incommensurable theories? Philosophy of Science, 46(2):244–262, June 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187047.

Mills:1979:PIF

[673] Susan K. Mills and John H. Beatty. The propensity interpretation of fitness. Philosophy of Science, 46(2):263–286, June 1979. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187048.

Wolfson:1979:SES

[674] Paul Wolfson and James Woodward. Scientific explanation and Sklar’s views of space and time. Philosophy of Science, 46(2):287–294, June 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187049.

Simon:1979:FFU

[675] Herbert A. Simon. Fit, finite, and universal axiomatization of theories. Philosophy of Science, 46(2):295–301, June 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187050.

Torretti:1979:JPC

[676] Roberto Torretti. Jackson and Pargetter’s criterion of distant simultaneity. Philosophy of Science, 46(2):302–305, June 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187051.

Giannoni:1979:CBR

[677] Carlo Giannoni. Comment on “Relative Simultaneity in the Special Theory of Relativity”. Philosophy of Science, 46(2):306–309, June 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187052. See [551].

Jackson:1979:RTG

[678] Frank Jackson and Robert Pargetter. A reply to Torretti and Giannoni. Philosophy of Science, 46(2):310–315, June 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187053.

Hull:1979:RG

[679] David L. Hull. Reduction in genetics. Philosophy of Science, 46(2):316– 320, June 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187054.

Clark:1979:TMM

[680] Desmond M. Clark. Teleology and mechanism: M. Grene’s absurdity argument. Philosophy of Science, 46(2):321–325, June 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187055. See comment [681].

Grene:1979:CDC

[681] Marjorie Grene. Comment on Desmond Clarke, “Teleology and Mechanism: M. Grene’s Absurdity Argument”. Philosophy of Science, 46(2): 326–327, June 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187056. See

[680].

Weingard:1979:GRC

[682] Robert Weingard. General Relativity and the conceivability of time travel. Philosophy of Science, 46(2):328–332, June 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187057.

Heimann:1979:BRB

[683] P. M. Heimann. Book review: Epistemological Writings by Hermann Von Helmholtz; Malcolm F. Lowe; Robert S. Cohen; Yehuda Elkana. Philosophy of Science, 46(2):333–334, June 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187058.

Mitroff:1979:BRB

[684] Ian I. Mitroff. Book review: Perspectives in the Sociology of Science by Stuart S. Blume. Philosophy of Science, 46(2):334–335, June 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187059.

Macklin:1979:BRB

[685] Ruth Macklin. Book review: Philosophical Medical Ethics: Its Nature and Significance by Stuart F. Spicker; H. Tristram Englehardt. Philosophy of Science, 46(2):335–336, June 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187060.

Demopoulos:1979:BRB

[686] William Demopoulos. Book review: The Uncertainty Principle and Foundations of Quantum Mechanics: A Fifty Years’ Survey by W. Price; S. Chissick. Philosophy of Science, 46(2):336–338, June 1979. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187061.

Curd:1979:BRB

[687] Martin V. Curd. Book review: Scientific Materialism in Nineteenth Century Germany by Frederick Gregory. Philosophy of Science, 46(2):338–340, June 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/187062.

Anonymous:1979:BMb

[688] Anonymous. Back matter. Philosophy of Science, 46(2):341, June 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187063.

Anonymous:1979:FMb

[689] Anonymous. Front matter. Philosophy of Science, 46(2):??, June 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187043.

Enc:1979:FAF

[690] Berent Enc. Function attributions and functional explanations. Philosophy of Science, 46(3):343–365, September 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186848.

Meixner:1979:HED

[691] John Meixner. Homogeneity and explanatory depth. Philosophy of Science, 46(3):366–381, September 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186849.

Shames:1979:MDE

[692] Morris L. Shames. On the metamethodological dimension of the “expectancy paradox”. Philosophy of Science, 46(3):382–388, September 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186850.

Kimbrough:1979:RGM

[693] Steven Orla Kimbrough. On the reduction of genetics to molecular biology. Philosophy of Science, 46(3):389–406, September 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186851.

Levin:1979:TCM

[694] Michael E. Levin. On theory-change and meaning-change. Philosophy of Science, 46(3):407–424, September 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186852.

Joseph:1979:GSR

[695] Geoffrey Joseph. Geometry and Special Relativity. Philosophy of Science, 46(3):425–438, September 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186853.

Nerlich:1979:CIP

[696] Graham Nerlich. Is curvature intrinsic to physical space? Philosophy of Science, 46(3):439–458, September 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186854.

Berkowitz:1979:AES

[697] Leonard J. Berkowitz. Achinstein on empirical significance: A matter of principle. Philosophy of Science, 46(3):459–465, September 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186855.

Jones:1979:CJI

[698] Gary Jones. Clendinnen, Jackson, and induction. Philosophy of Science, 46(3):466–469, September 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 186856. See [49, 48].

Hooker:1979:BRH

[699] C. A. Hooker. Book review: Hollis and Nell’s Rational Economic Man: A Philosophical Critique of Neo-Classical Economics. Philosophy of Science, 46(3):470–490, September 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186857.

Giere:1979:BRB

[700] Ronald N. Giere. Book review: Inference, Method and Decision by Roger D. Rosenkrantz. Philosophy of Science, 46(3):491–493, September 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186858.

Salmon:1979:BRB

[701] Marrilee H. Salmon. Book review: Time and Traditions: Essays in Archaeological Interpretation by Bruce Trigger. Philosophy of Science, 46 (3):494–495, September 1979. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186859.

Osler:1979:BRB

[702] Margaret J. Osler. Book review: New Perspectives on Galileo by R. E. Butts; J. C. Pitt. Philosophy of Science, 46(3):495–496, September 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186860.

McEvoy:1979:BRB

[703] John G. McEvoy. Book review: Philosophy and Methodology in the Social Sciences by Barry Hindess. Philosophy of Science, 46(3):496–499, September 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186861.

Anonymous:1979:BMc

[704] Anonymous. Back matter. Philosophy of Science, 46(3):??, September 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186862.

Anonymous:1979:FMc

[705] Anonymous. Front matter. Philosophy of Science, 46(3):??, September 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186847.

MacKinnon:1979:SRN

[706] Edward MacKinnon. Scientific realism: The new debates. Philosophy of Science, 46(4):501–532, December 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187247.

Richardson:1979:FR

[707] Robert C. Richardson. Functionalism and reductionism. Philosophy of Science, 46(4):533–558, December 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187248.

Green:1979:GO

[708] Michael B. Green. The grain objection. Philosophy of Science, 46(4): 559–589, December 1979. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187249.

Shope:1979:EMA

[709] Robert K. Shope. Eliminating mistakes about eliminative materialism. Philosophy of Science, 46(4):590–612, December 1979. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187250.

Kitts:1979:BSN

[710] David B. Kitts and David J. Kitts. Biological species as natural kinds. Philosophy of Science, 46(4):613–622, December 1979. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187251.

McMullin:1979:BRL

[711] Ernan McMullin. Book review: Laudan’s Progress and Its Problems. Philosophy of Science, 46(4):623–644, December 1979. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187252.

Cupples:1979:MSE

[712] Brian Cupples. Moor and Schlesinger on explanation. Philosophy of Science, 46(4):645–650, December 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187253.

Sayre:1979:BRB

[713] Kenneth M. Sayre. Book review: The Computer Revolution in Philosophy: Philosophy, Science and Models of Mind by Aaron Sloman. Philosophy of Science, 46(4):651–652, December 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187254.

Lennox:1979:BRB

[714] James G. Lennox. Book review: Explanatory Structures: A Study of Concepts of Explanation in Early Physics and Philosophy by Stephen Gaukroger. Philosophy of Science, 46(4):652–654, December 1979. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187255.

Nagel:1979:BRB

[715] Gordon Nagel. Book review: Kant’s Theory of Science by Gordon G. Brittan. Philosophy of Science, 46(4):654–655, December 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187256.

Strong:1979:BRB

[716] John V. Strong. Book review: Peirce’s Philosophy of Science. Critical Studies in His Theory of Induction and Scientific Method by Nicholas Rescher. Philosophy of Science, 46(4):655–657, December 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187257.

Morton:1979:BRB

[717] Adam Morton. Book review: Studies in Perception by Peter K. Machamer; Robert G. Turnbull. Philosophy of Science, 46(4):657–659, December 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187258.

Anonymous:1979:MRR

[718] Anonymous. In memorium: Richard S. Rudner 1921–1979. Philosophy of Science, 46(4):660, December 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187259.

Anonymous:1979:BMd

[719] Anonymous. Back matter. Philosophy of Science, 46(4):661–664, December 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187260.

Anonymous:1979:VI

[720] Anonymous. Volume information. Philosophy of Science, 46(4):665–669, December 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187245.

Anonymous:1979:FMd

[721] Anonymous. Front matter. Philosophy of Science, 46(4):??, December 1979. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187246.